### v1

Abstract.
In this study, we investigate the incompressible generalised Navier-Stokes-Voigt equations within a bounded domain $\Omega\subset\mathbb{R}^{d}$, where $d\geq 2$. The governing momentum equation is expressed as:
$\partial_{t}(\bm{v}-\kappa\Delta\bm{v})+\nabla\cdot(\bm{v}\otimes\bm{v})+\nabla\pi-\nu\nabla\cdot\mathopen{}\mathclose{{\left(|\mathbf{D}(\bm{v})|^{p-2}\mathbf{D}(\bm{v})}}\right)=\bm{f}.$
Here, for $d\in\{2,3\}$, $\bm{v}$ represents the velocity field, $\pi$ denotes the pressure, and $\bm{f}$ is the external forcing term. The constants $\kappa$ and $\nu$ correspond to the relaxation time and kinematic viscosity, respectively. The parameter $p\in(1,\infty)$ characterizes the fluid’s flow behavior, and $\mathbf{D}(\bm{v})$ denotes the symmetric part of the velocity gradient $\nabla\bm{v}$. For the power-law exponent $p\in\mathopen{}\mathclose{{\left(\frac{2d}{d+2},\infty}}\right)$, we establish the existence of a weak solution to the generalised Navier-Stokes-Voigt equations. Furthermore, we demonstrate that the weak solution is unique for the same range of the exponent $p$. The optimality of our results lies in the framework’s use of a Gelfand triple, which allows the Aubin–Dubinskiĭ lemma to yield strong convergence of approximate solutions, essential for existence and valid precisely for $p>\frac{2d}{d+2}$.
1. Introduction
1.1. The model
We study flows of incompressible fluids and constant density with elastic properties governed by the incompressible generalised Navier-Stokes-Voigt (NSV) equations in a bounded domain $\Omega\subset\mathbb{R}^{d}$, $d\geq 2$, with its boundary denoted by $\partial\Omega$, during the time interval $[0,T]$, with $0<T<\infty$. The momentum equation is assumed here to be perturbed by a relaxation term $\kappa\Delta{\partial_{t}\bm{v}}$ which makes the problem posed by the following system of equations:
[…]
where $Q_{T}:=\Omega\times(0,T)$ and $\Gamma_{T}:=\partial\Omega\times(0,T)$. Here, $\bm{v}=(v_{1},\dots,v_{d})$ and $\bm{f}=(f_{1},\dots,f_{d})$ are vector-valued functions, $\pi$ is a scalar-valued function, whereas $\nu$ and $\kappa$ are positive constants. In real-world applications, space dimensions of interest are $d=2$ and $d=3$, and in that case $\bm{v}$ accounts for the velocity field, $\pi$ is the pressure (divided the constant density), $\bm{f}$ represents an external forces field. Greek letters $\nu$ and $\kappa$ stay for positive constants that, in the dimensions of physics interest, account for the kinematics viscosity and for the relaxation time, that is the time required for a viscoelastic fluid relax from a deformed state back to its equilibrium configuration. The power-law exponent $p$ is a constant that characterizes the flow and is assumed to be such that $1<p<\infty$. Capital letters stay for tensor-valued functions, in particular $\mathbf{I}$ is the unit tensor and $\mathbf{D}(\bm{v})$ is the symmetric part of the velocity gradient, i.e. $\mathbf{D}(\bm{v}):=\frac{1}{2}\mathopen{}\mathclose{{\left(\nabla\bm{v}+(\nabla\bm{v})^{\top}}}\right)$. The notation $\partial_{t}\bm{v}$} stays, as usual, for the partial time derivative of $\bm{v}$.
Equation (1.1)1 can be derived from the principle of conservation of momentum by considering the deviatoric part of the Cauchy stress tensor of the form
$\bm{\tau}=2\nu\,|\mathbf{D}(\bm{v})|^{p-2}\,\mathbf{D}(\bm{v})+2\kappa\,\partial_{t}\mathbf{D}(\bm{v}),$ (1.2)
which represents a nonlinear viscous response with an additional linear elastic contribution. When $\kappa=0$, this extra stress reduces to the pure power-law (Ostwald–de Waele) relation
$\bm{\tau}=2\nu\bigl(|\mathbf{D}(\bm{v})|\bigr)\,\mathbf{D}(\bm{v}),\qquad\nu(s)=\nu_{0}\,s^{p-2}\quad(p>1,\ \nu_{0}>0),$ (1.3)
so that the apparent viscosity scales like $|\mathbf{D}(\bm{v})|^{p-2}$. In this setting, shear-thinning (or pseudo-plastic) fluids correspond to exponents $1<p<2$, shear-thickening (or dilatant) fluids to $p>2$, whereas the special case $p=2$ recovers the classical Navier-Stokes equations for Newtonian fluids. When $p=2$ and $\kappa>0$ (with $\nu>0$) in the constitutive law (1.2), the deviatoric stress tensor reflects a combined effect of Newton’s law of viscosity and Hooke’s law of linear elasticity. For this case, equation (1.2) generalizes the one-dimensional Kelvin-Voigt model to higher dimensions and has been successfully applied to describe the rheology of viscous fluids with elastic properties, such as polymer solutions and certain biological tissues [46, 17]. The classical one-dimensional Kelvin-Voigt model,
$\tau=\nu\,\frac{\partial v}{\partial x}+\kappa\,\frac{\partial^{2}v}{\partial x\,\partial t},$ (1.4)
provides a simple constitutive relation for linear viscoelastic materials. This model can be obtained either from the Boltzmann superposition principle or through mechanical analogs consisting of elastic springs and viscous dashpots [9]. A natural power-law generalization of (1.4), accounting for non-Newtonian viscous fluids endowed with (linear) elastic properties, is given by
$\tau=\nu\,\mathopen{}\mathclose{{\left|\frac{\partial v}{\partial x}}}\right|^{p-2}\,\frac{\partial v}{\partial x}+\kappa\,\frac{\partial^{2}v}{\partial x\,\partial t}.$ (1.5)
In fluid mechanics, the choice of a constitutive equation is largely directed by the specific application under consideration. For describing the rheological behavior of many incompressible viscoelastic fluids, equations of the type (1.4) and (1.5) are widely regarded as a natural starting point [12]. Since a great variety of stress–strain relations may be postulated, it is customary to impose admissibility criteria in order to restrict attention to physically meaningful models. For incompressible fluids, Oldroyd [44] proposed that constitutive laws should be: (a) frame-invariant, (b) deterministic, and (c) local. However, the elastic tensor $\partial_{t}\mathbf{D}(\bm{v})$ fails to be objective and thus does not satisfy the principle of frame indifference. To illustrate this, let $\mathbf{D}(\bm{v})^{\ast}$ and $\partial_{t}\mathbf{D}(\bm{v})^{\ast}$ denote the representations of $\mathbf{D}(\bm{v})$ and $\partial_{t}\mathbf{D}(\bm{v})$ in another reference frame, and let $\mathbf{Q}=\mathbf{Q}(t)$ be an arbitrary orthogonal tensor. Since $\mathbf{D}(\bm{v})$ is objective, we have $\mathbf{D}(\bm{v})^{\ast}=\mathbf{Q}\mathbf{D}(\bm{v})\mathbf{Q}^{\top}$. Objectivity of $\partial_{t}\mathbf{D}(\bm{v})$ would then require
$\partial_{t}\mathbf{D}(\bm{v})^{\ast}=\partial_{t}\mathbf{Q}\mathbf{D}(\bm{v})\mathbf{Q}^{\top}+\mathbf{Q}\partial_{t}\mathbf{D}(\bm{v})\mathbf{Q}^{\top}+\mathbf{Q}\mathbf{D}(\bm{v})\partial_{t}\mathbf{Q}^{\top},$
so that
$\partial_{t}\mathbf{Q}\mathbf{D}(\bm{v})\mathbf{Q}^{\top}+\mathbf{Q}\mathbf{D}(\bm{v})\partial_{t}\mathbf{Q}^{\top}=\bm{0},$
a condition not generally satisfied. To obtain an admissible model, the partial time derivative in (1.3) can be replaced by the convected time derivative [12]
$\mathbf{D}(\bm{v})^{\triangledown}=\partial_{t}\mathbf{D}(\bm{v})+(\bm{v}\cdot\nabla)\mathbf{D}(\bm{v})-\Big(\mathbf{D}(\bm{v})\nabla\bm{v}+(\mathbf{D}(\bm{v})\nabla\bm{v})^{\top}\Big).$ (1.6)
However, the PDEs obtained from constitutive relations involving (1.6) are highly nonlinear and analytically difficult, so certain simplifications are required to make a rigorous mathematical study feasible.
1.2. Literature review
The mathematical analysis of pure ($p\not=2$) power-law fluids was initiated by Ladyzhenskaya [33, 34], followed by Lions [40]. In her celebrated papers [33, 34], Ladyzhenskaya proposed studying what she termed the modified Navier–Stokes equations, in which the deviatoric stress tensor is prescribed by (1.3). Models of this kind had already been investigated experimentally since the early twentieth century in connection with polymer melts, colloidal suspensions, and other non-Newtonian fluids (see e.g. [9]). By adopting this form of the stress–strain relation, Ladyzhenskaya naturally placed her mathematical framework in line with experimentally observed non-Newtonian behavior, even though her main goal was to establish solvability results rather than to develop a detailed physical theory. In [33, 34], it was proved the existence of 3d weak solutions for
$p\geq\frac{11}{5},$
and their uniqueness for
$p\geq\frac{5}{2},$
Subsequently, Lions [40] extended these results to the case of a general space dimension $d\geq 2$, and where the diffusion depends on the full velocity gradient $\nabla\bm{v}$ in place of its symmetric part $\mathbf{D}(\bm{v})$. He established the existence for
$p\geq\frac{3d+2}{d+2},$ (1.7)
and uniqueness for
$p\geq\frac{d+2}{2}.$
The existence results in [33, 34, 40], in addition to employing Galerkin approximations combined with compactness argumnets and the theory of monotone operators, rely (directly or indirectly) on the continuous embedding
$\mathrm{L}^{p}(0,T;\mathbf{V}_{p})\cap\mathrm{L}^{\infty}(0,T;\mathbf{H})\hookrightarrow\mathrm{L}^{\frac{p(d+2)}{d}}(0,T;\mathrm{L}^{\frac{p(d+2)}{d}}(\Omega)^{d}),$ (1.8)
which ensures that $\bm{v}\otimes\bm{v}:\mathbf{D}(\bm{\varphi})$ is bounded in $\mathbf{L}^{1}(Q_{T})$ for all $\bm{v}$ and $\bm{\varphi}$ in the same function spaces whenever (1.7) holds. See the next section for the definition of the function spaces appearing in (1.8).
More than forty years later, Zhikov [54] succeeded in improving the existence results [33, 34, 40] to the case of
$p\geq\max\mathopen{}\mathclose{{\left\{\frac{3d}{d+2},\frac{d+\sqrt{3d^{2}+4d}}{d+2}}}\right\}.$
The main idea of his proof consisted in a regularization of the deviatoric stress tensor followed by the application of the classical existence results [33, 34, 40]. The passage to the limit in the regularized problem is justified using arguments from Measure Theory, which allow one to handle the weak convergence of nonlinear terms. It should be stressed that this lower bound for $p$ had already been obtained by Nečas and his collaborators [41], although their analysis was carried out in the setting of periodic boundary conditions instead.
Even prior to Zhikov’s result, Wolf [51] established the existence of weak solutions to the Navier-Stokes system with a power-law viscosity for exponents
$p>\frac{2(d+1)}{d+2}.$
In his proof, Wolf employed a decomposition of the pressure into a regular (measurable) part and a singular part, which allowed him to handle the nonlinear convective term effectively. This was combined with the $\mathrm{L}^{\infty}$–truncation method, a technique that enables control of the nonlinearities and provides compactness necessary to pass to the limit in the weak formulation.
Finally, Diening et al. [21] extended and improved the existence result [51] to cover smaller values of the power-law exponent $p$, specifically
$\frac{2d}{d+2}<p<2.$
In this work, the main tool was the successful combination of the pressure decomposition technique [51] with the Lipschitz–truncation method, which provides strong control over approximate solutions. This approach enabled the authors to extend the existence theory of weak solutions to a wider range of power-law exponents, particularly in the shear-thinning case.
Concerning existence, the range
$1<p<\frac{2d}{d+2},\quad\mbox{for}\ \ d\not=2,$
remains out of reach with the techniques currently available. The main difficulty lies in the lack of an appropriate Gelfand triple that would allow the application of the Aubin-Dubinskiĭ compactness lemma to extract strong convergence of approximate solutions.
With regard to uniqueness for the Navier-Stokes system with a power-law viscosity, the classical results of [33, 34, 40] have only been improved (for $d\geq 3$) up to the range
$p\geq\frac{3d+2}{\,d+2\,}.$
In particular, see the work by Du and Gunzburger [22] as well as the contributions by Nečas and his collaborators [41], where refined energy methods were used to establish uniqueness of weak (or suitably regular) solutions.
The particular case of the momentum (1.1)1 with $p=2$ is usually referred to as the Navier-Stokes-Voigt (NSV) equations. Over the past several decades, the corresponding problem and related models have attracted considerable attention in the mathematical literature due to their analytical tractability and relevance to applications. The first rigorous study of existence and uniqueness of weak solutions for the NSV equations is due to Oskolkov [45], who coined the name Kelvin-Voigt (initially misspelled Kelvin-Voight) for this type of regularized fluid model. As noted by Zvyagin and Turbin [55], neither Kelvin nor Voigt themselves proposed a constitutive stress–strain law or a complete set of governing equations for viscoelastic fluids. Nevertheless, the name Navier-Stokes-Voigt has become widely accepted, since this model naturally extends the equations originally suggested by Voigt (for elastic media with relaxation effects) to encompass viscous fluids with viscoelastic characteristics. From the viewpoint of applications, one of the earliest uses of this model appears to be due to Pavlovsky [46], who employed a simplified form of the equations to describe weakly concentrated water–polymer mixtures. More recently, Ebrahimi et al. [23] explored the two-dimensional NSV equations as a tool in numerical algorithms, including applications to image inpainting, illustrating the flexibility of the model beyond classical fluid dynamics. Since the pioneering contributions of Oskolkov [45], the NSV equations have been investigated extensively with respect to existence, uniqueness, regularity, and asymptotic properties of solutions. A key analytical feature, first emphasized by Ladyzhenskaya [32] and later by Oskolkov himself, is that the additional Voigt term $\kappa\partial_{t}\mathbf{D}(\bm{v})$ acts as a regularizing mechanism for the Navier-Stokes system: the three-dimensional transient problem possesses a unique global weak solution. Before the year 2000, most of the research on this model was carried out by the Russian School of Analysis (see the survey [55] and references therein). In the following years, the NSV equations have been further examined by many authors in various directions. Cao et al. [15] proposed the system as a physically motivated regularization (for small $\kappa$) of the three-dimensional Navier-Stokes equations for purposes of direct numerical simulation, both under periodic and no-slip Dirichlet boundary conditions. Several works have addressed the long-time dynamics of solutions: Titi and his collaborators [30, 29], and Coti Zelati and Gal [16] established existence and regularity of global attractors, while Dou and Qin [47] and Garcia-Luengo et al. [26] analyzed uniform and pullback attractors. Berselli and Bisconti [10] analyzed NSV structural stability under viscosity and relaxation variations, while Berselli and Spirito [11] constructed suitable weak solutions for the 3D Navier-Stokes equations via the Voigt approximation. The decay rates of solutions in relation to the initial data have been characterized by Niche [43] and Zhao and Zhu [52]. Connections to turbulence modeling have also been highlighted: Titi and his collaborators [36, 37] related the NSV equations to the Bardina turbulence models, and similar observations were made by Lewandowski et al. [38, 39]. Damázio et al. [18] establish an $\mathrm{L}^{q}$–theory for the NSV equations in bounded domains, proving existence, uniqueness, and regularity of solutions, while Baranovskii [8] analyzes the NSV equations with position-dependent slip boundary conditions, addressing well-posedness and the impact of such boundaries. Mohan [42] provided a comprehensive study of global solvability, uniqueness, asymptotic behavior, and certain control problems for the three-dimensional Navier-Stokes-Voigt equations with memory effects. Antontsev et al. [4, 6, 5] studied NSV equations with variable density, proving existence, uniqueness, and regularity, including cases with unbounded density, and extended the analysis to long-time behavior, anisotropic viscosity, relaxation, and damping. De Oliveira et al. [19] established the existence and uniqueness of strong solutions for the NSV equations with non-negative (possibly vanishing) density, providing detailed regularity and well-posedness results. Zvyagin and Turbin [56, 57] also established weak solvability for NSV equations with non-constant density, including finite-order models, even when the initial density is not strictly positive.
The power-law case, with the exponent $p$ varying over the entire interval $(1,\infty)$, has been investigated by relatively few authors to date. Antontsev et al. [2, 3, 7] studied the Navier-Stokes-Voigt equations incorporating a pp–Laplacian or anisotropic diffusion (allowing the power-law exponent to differ along distinct spatial directions), as well as relaxation and damping terms, and provided a detailed analysis of existence, uniqueness, conditions for finite-time blow-up, and the large-time behavior of solutions. In a stochastic setting, Kumar et al. [31] established existence and uniqueness of weak solutions for the generalised (power-law) stochastic Navier-Stokes-Voigt equations, thereby providing a rigorous mathematical framework for studying this model under random forcing and highlighting the impact of stochasticity on the solution dynamics.
1.3. Objectives and novelties of the paper
The main goal of this article is to demonstrate the existence of weak solutions to the generalised Navier–Stokes–Voigt equations. The literature on this model appears to be limited, and to the best of our knowledge, the analysis presented here has not yet been explored.
  • As mentioned above, there are few works in the literature addressing power-law fluids [2, 3, 7, 21, 51]. In [51], the authors studied power-law fluids and established existence results for the exponent $p\in\big(\frac{2(d+1)}{d+2},\infty\big)$ using the $\mathrm{L}^{\infty}$-truncation, while [21] extended these results to $p\in\big(\frac{2d}{d+2},\infty\big)$ via Lipschitz-truncation. In our analysis, the presence of the Voigt term $\kappa\partial_{t}\Delta\bm{v}$ allows us to establish existence and uniqueness of the generalised Navier–Stokes–Voigt equations for the optimal range $p\in\big(\frac{2d}{d+2},\infty\big)$ using compactness and monotonicity arguments.
  • Another novelty of this article is that we do not employ the Helmholtz–Hodge projection, which would restrict our analysis to the case $p\in[2,\infty)$. Although the Helmholtz–Hodge projection is well-defined for $p\in(1,2)$, the analysis in this range is highly delicate and creates significant difficulties in controlling the convective term in the momentum equation (1.1)1. By avoiding the Helmholtz–Hodge projection, we are able to treat the full range $p>\frac{2d}{d+2}$ directly, without encountering these technical obstacles.
1.4. Organization
The structure of this article is organized as follows: Section 2 lays the groundwork by introducing the necessary function spaces and auxiliary results essential for the analysis. In Section 3, we present the definition of the weak solution (Definition 3.1) and state our main results, including Theorems 3.2 and 3.3. As previously mentioned, the proof of Theorem 3.2 is divided into two cases based on the range of the exponent $p$ (see (3.6)). Section 4 focuses on the proof of Theorem 3.2 for the case of large $p$, specifically when $p\geq\frac{3d-4}{d}$. This section leverages key tools such as the Banach-Alaoglu theorem (Subsection 4.3) and the Minty trick (Subsection 4.4). In Section 5, we address the decomposition of the pressure term, ensuring that each decomposed component corresponds to a distinct equation, as formalized in Theorem 5.1. The proof of Theorem 3.2 for small values of $p$, where $p\in\mathopen{}\mathclose{{\left(\frac{2d}{d+2},\frac{3d-4}{d}}}\right)$, is carried out in Section 6. This section introduces the regularized problem (6.3) and its corresponding definition (Definition 6.1). It also reconstructs the pressure term (Subsection 6.3) and employs compactness arguments (Subsection 6.4) alongside the monotonicity method (Subsection 6.5). Finally, the article concludes with Section 7, where we establish the uniqueness of the weak solution, as stated in Theorem 3.3.
2. Auxiliary results
2.1. Function spaces
Let us introduce the functions spaces used to characterize the solutions of Fluid Mechanics problems. Let $\Omega$ be a domain in $\mathbb{R}^{d}$, that is, an open and connected subset of $\mathbb{R}^{d}$. By $\mathrm{C}_{0}^{\infty}(\Omega)^{d}$ we denote the space of all infinitely differentiable $\mathbb{R}^{d}$-valued functions with compact support in $\Omega$. For $1\leq p\leq\infty$, we denote by $\mathrm{L}^{p}(\Omega)^{d}$ the Lebesgue space consisting of all $\mathbb{R}^{d}$-valued measurable (equivalent classes of) functions that are $p$-summable over $\Omega$. By $|\Omega|$ we denote the dd–Lebesgue measure of $\Omega$, and $\lambda^{d}$ accounts for the $d$-dimensional Lebesgue measure. The corresponding Sobolev spaces are represented by $\mathrm{W}^{k,p}(\Omega)^{d}$, where $k\in\mathbb{N}$. When $p=2$, $\mathrm{W}^{k,2}(\Omega)^{d}$ are Hilbert spaces that we denote by $\mathrm{H}^{k}(\Omega)^{d}$. We define
$\bm{\mathcal{V}}$ $:=\big\{\bm{v}\in\mathrm{C}_{0}^{\infty}(\Omega)^{d}:\nabla\cdot\bm{v}=0\big\},$
[…]
for $1\leq p<\infty$ and $k\in\mathbb{N}$. In the particular case of $p=2$, we denote the spaces $\mathbf{L}_{\sigma}^{p}(\Omega)$ and $\mathbf{V}_{p}^{k}$ by $\mathbf{H}$ and $\mathbf{V}^{k}$, respectively. If $k=1$, we denote $\mathbf{V}_{p}^{k}$ by $\mathbf{V}_{p}$, and if both $p=2$ and $k=1$, we denote $\mathbf{V}_{p}^{k}$ solely by $\mathbf{V}$. Let $(\cdot,\cdot)$ stand for the inner product of the Hilbert space $\mathrm{L}^{2}(\Omega)^{d}$, and we denote by $\langle\cdot,\cdot\rangle,$ the induced duality product between $\mathrm{W}_{0}^{1,p}(\Omega)^{d}$ and its dual space $\mathrm{W}^{-1,p^{\prime}}(\Omega)^{d}$, as well as between $\mathrm{L}^{p}(\Omega)^{d}$ and its dual $\mathrm{L}^{p^{\prime}}(\Omega)^{d}$, where $\frac{1}{p}+\frac{1}{p^{\prime}}=1$. In the sequel, the $\mathrm{L}^{p}$, $\mathrm{W}^{k,p}$ and $\mathrm{W}^{-k,p^{\prime}}$ norms will be denoted in short by $\|\cdot\|_{p}$, $\|\cdot\|_{k,p}$ and $\|\cdot\|_{-k,p^{\prime}}$. On $\mathbf{V}$, we consider the equivalent norm $\|\bm{v}\|_{\mathbf{V}}:=\|\nabla\bm{v}\|_{2}$, $\bm{v}\in\mathbf{V}$.
Given a Banach space $\mathbf{B}$, we denote by $\mathrm{L}^{r}(0,T;\mathbf{B})$, with $1\leq r\leq\infty$, the space consisting of all Bochner $\mathbf{B}$-valued measurable (equivalent classes of) functions that are $r$-summable over $[0,T]$. In particular, $\mathrm{C}_{\rm{w}}([0,T];\mathbf{B})$ denotes the subspace of $\mathrm{L}^{\infty}(0,T;\mathbf{B})$ formed by weakly continuous functions from $[0,T]$ into $\mathbf{B}$. We also consider the following function space
$\mathbf{C}^{\infty}_{0,\operatorname{div}}(Q_{T}):=\mathopen{}\mathclose{{\left\{\bm{\varphi}\in\mathrm{C}^{\infty}(Q_{T})^{d}:\mathbf{div}\,\bm{\varphi}=0\ \text{ with }\operatorname{supp}(\bm{\varphi})\Subset\Omega\times[0,T)}}\right\}.$
2.2. Auxiliary results
We recall the following inequalities which are classical in the theory of $p$-Laplace equations.
Lemma 2.1.
For all $\mathbf{E},\ \mathbf{F}\in\mathbb{R}^{d\times d}$, the following assertions hold true:
$\frac{1}{2^{p-1}}|\mathbf{E}-\mathbf{F}|^{p}$ $\leq\big(|\mathbf{E}|^{p-2}\mathbf{E}-|\mathbf{F}|^{p-2}\mathbf{F}\big):\big(\mathbf{E}-\mathbf{F}\big),$ $\text{ for }p\in[2,\infty),$ (2.1)
$(p-1)|\mathbf{E}-\mathbf{F}|^{2}$ $\leq\big(|\mathbf{E}|^{p-2}\mathbf{E}-|\mathbf{F}|^{p-2}\mathbf{F}\big):(\mathbf{E}-\mathbf{F})\big(|\mathbf{E}|^{p}+|\mathbf{F}|^{p}\big)^{\frac{2-p}{p}},$ $\text{ for }p\in(1,2).$ (2.2)
Proof.
The proof combines Lemmas 5.1 and 5.2 of [28]. ∎
From (2.1) and (2.2), one easily gets
$\big(|\mathbf{D}(\bm{v})|^{p-2}\mathbf{D}(\bm{v})-|\mathbf{D}(\bm{u})|^{p-2}\mathbf{D}(\bm{u})\big):\big(\mathbf{D}(\bm{v})-\mathbf{D}(\bm{u})\big)\geq 0,\ \text{ for all }\ \bm{v},\bm{u}\in\mathbf{V}_{p},$ (2.3)
and whenever $1<p<\infty$. As a consequence of (2.3), the operator
$\mathbf{A}(\bm{v}):=|\mathbf{D}(\bm{v})|^{p-2}\mathbf{D}(\bm{v})$ (2.4)
is said to be monotone for any $p$ such that $1<p<\infty$.
The following result generalises Simon [48, Corollaries 8 and 9] versions of Aubin-Dubinskiĭ compactness lemma. Right below, we assume $p_{0},\ p_{1}\in[1,\infty]$, $s_{0},\ s_{1}\in\mathbb{R}$ and $T>0$, denote $\partial_{t}x$ as the distributional time derivative of $x$, and define $p_{\theta}$ and $s_{\theta}$ as follows
$\frac{1}{p_{\theta}}:=\frac{1-\theta}{p_{0}}+\frac{\theta}{p_{1}},\quad s_{\theta}:=(1-\theta)s_{0}+\theta s_{1},\quad\theta\in(0,1).$
Lemma 2.2.
Let $\mathrm{E}$, $\mathrm{E}_{0}$ and $\mathrm{E}_{1}$ be Banach spaces such that
$\mathrm{E}_{1}\hookrightarrow\hookrightarrow\mathrm{E}\hookrightarrow\mathrm{E}_{0}\quad\text{ and }\quad\|x\|_{\mathrm{E}}\leq C\|x\|_{\mathrm{E}_{0}}^{1-\theta}\|x\|_{\mathrm{E}_{1}}^{\theta},\text{ for any }\ x\in\mathrm{E}_{1}.$ (2.5)
  1. (1)
    Then
    $\mathrm{W}^{s_{0},p_{0}}(0,T;\mathrm{E}_{0})\cap\mathrm{W}^{s_{1},p_{1}}(0,T;\mathrm{E}_{1})\hookrightarrow\hookrightarrow\mathrm{W}^{s,p}(0,T;\mathrm{E}),\quad\mbox{if}\quad s<s_{\theta}\text{ and }s-\frac{d}{p}<s_{\theta}-\frac{d}{p_{\theta}};$
    $\mathrm{W}^{s_{0},p_{0}}(0,T;\mathrm{E}_{0})\cap\mathrm{W}^{s_{1},p_{1}}(0,T;\mathrm{E}_{1})\hookrightarrow\hookrightarrow\mathrm{C}^{s}([0,T];\mathrm{E}),\quad\mbox{if}\quad 0\leq s<s_{\theta}-\frac{d}{p_{\theta}}.$
  2. (2)
    If, in addition to (2.5),
    $\mbox{$\mathrm{X}$ is a bounded subset of $\mathrm{L}^{p_{1}}(0,T;\mathrm{E}_{1})$},$
    $\mbox{$\partial\mathrm{X}:=\mathopen{}\mathclose{{\left\{\partial_{t}x:x\in\mathrm{X}}}\right\}$ is bounded in $\mathrm{L}^{p_{0}}(0,T:\mathrm{E}_{0})$},$
    then $\mathrm{X}$ is relatively compact in:
    $\mathrm{W}^{s,p}(0,T;\mathrm{E}),\quad\mbox{if}\quad 0\leq s<1-\theta\text{ and }s-\frac{1}{p}<1-\theta-\frac{1}{p_{\theta}};$
    $\mathrm{C}^{s}([0,T];\mathrm{E}),\quad\mbox{if}\quad 0\leq s<1-\theta-\frac{1}{p_{\theta}}.$
Proof.
For the proof of (1), we address the reader to [1, Theorem 5.2], and for (2) to [1, Theorem 1.1]. See also the proofs of [48, Corollaries 8 and 9]. ∎
The following result provides a variant of the classical de Rham theorem.
Lemma 2.3.
Assume $1<\gamma<\infty$. For each $\bm{u}^{\ast}\in\mathrm{W}^{-1,\gamma}(\Omega)^{d}$ such that
$\big<\bm{u}^{\ast},\bm{u}\big>=0,\quad\text{ for all }\ \bm{u}\in\mathbf{V}_{\gamma^{\prime}},$
there exists a unique $\pi\in\mathrm{L}^{\gamma}(\Omega)$, with $\int_{\Omega}\pi\,\,\mathrm{d}\bm{x}=0$, such that
$\mathopen{}\mathclose{{\left\langle\bm{u}^{\ast},\bm{u}}}\right\rangle=\int_{\Omega}\pi\mathbf{div}\,\bm{u}\,\,\mathrm{d}\bm{x},\quad\text{ for all }\,\bm{u}\in\mathrm{W}_{0}^{1,\gamma^{\prime}}(\Omega)^{d}.$
Moreover, there exists a positive constant $C$ such that
$\|\pi\|_{\gamma}\leq C\|\bm{u}^{\ast}\|_{-1,\gamma}.$
Proof.
The proof is due to Bogovskiǐ [13] (see also [25, Theorems III.3.1 and III.5.3] and [50, Proposition I.1.1]). ∎
3. Main results
We now introduce the notion of weak solutions to the problem (1.1), which is the focus of our study. At this stage, we impose the following minimal assumptions on the problem data to guarantee that the subsequent definition is meaningful,
$\bm{f}\in\mathbf{L}^{1}(Q_{T}),\qquad\bm{v}_{0}\in\mathbf{V},\qquad\frac{2d}{d+2}\leq p<\infty.$ (3.1)
Definition 3.1.
Assume that conditions (3.1) are satisfied. We say that $\bm{v}$ is a weak solution to the problem (1.1), if:
  1. (1)
    $\bm{v}\in\mathrm{L}^{\infty}(0,T;\mathbf{V})\cap\mathrm{L}^{p}(0,T;\mathbf{V}_{p})$;
  2. (2)
    ${\partial_{t}\bm{v}}\in\mathrm{L}^{2}(0,T;\mathbf{V})$;
  3. (3)
    $\bm{v}(0)=\bm{v}_{0}$;
  4. (4)
    The following identity,
    $\begin{split}&\hskip-5.69046pt-\!\int_{0}^{T}\!\!\!\!\int_{\Omega}\bm{v}\cdot\partial_{t}\bm{\varphi}\,\mathrm{d}\bm{x}\mathrm{d}t-\kappa\!\int_{0}^{T}\!\!\!\!\int_{\Omega}\nabla\bm{v}:\nabla\partial_{t}\bm{\varphi}\,\mathrm{d}\bm{x}\mathrm{d}t+\nu\!\int_{0}^{T}\!\!\!\!\int_{\Omega}|\mathbf{D}(\bm{v})|^{p-2}\mathbf{D}(\bm{v}):\mathbf{D}(\bm{\varphi})\,\mathrm{d}\bm{x}\mathrm{d}t\\ &\hskip-5.69046pt=\int_{\Omega}\bm{v}_{0}\cdot\bm{\varphi}(0)\,\mathrm{d}\bm{x}+\kappa\int_{\Omega}\nabla\bm{v}_{0}:\nabla\bm{\varphi}(0)\,\mathrm{d}\bm{x}+\!\int_{0}^{T}\!\!\!\!\int_{\Omega}\bm{f}\cdot\bm{\varphi}\,\mathrm{d}\bm{x}\mathrm{d}t+\!\int_{0}^{T}\!\!\!\!\int_{\Omega}\bm{v}\otimes\bm{v}:\mathds{\nabla}\bm{\varphi}\,\mathrm{d}\bm{x}\mathrm{d}t,\end{split}$ (3.2)
    holds for all $\bm{\varphi}\in\mathbf{C}^{\infty}_{\operatorname{div}}(Q_{T})$ with $\operatorname{supp}(\bm{\varphi})\Subset\Omega\times[0,T)$.
As usual, condition $\bm{v}(0)=\bm{v}_{0}$ is interpreted in the sense of weak continuity in time, that is
$\lim_{t\to 0^{+}}\int_{\Omega}\bm{v}(t)\cdot\bm{\psi}=\int_{\Omega}\bm{v}_{0}\cdot\bm{\psi},\quad\text{ for all }\ \bm{\psi}\in\bm{\mathcal{V}}.$
The main result of this work concerns the existence of a weak solution to problem (1.1). To prove this result, we strengthen the conditions in (3.1) on the forcing term, initial velocity, and range of the power-law exponent. In addition to the condition (3.5) imposed below, which strengthens the admissible range of the power-law exponent, we assume
$\bm{v}_{0}\in\mathbf{V}\cap\mathbf{V}_{p},$ (3.3)
${\bm{f}\in\mathrm{L}^{p^{\prime}}(0,T;\mathrm{L}^{p^{\prime}}(\Omega)^{d})}.$ (3.4)
Theorem 3.2.
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded domain with a smooth boundary $\partial\Omega$ of class $\mathrm{C}^{2}$, and assume that hypotheses (3.3)-(3.4) are verified. If $2\leq d\leq 4$ and
$p>\frac{2d}{d+2},$ (3.5)
then there exists, at least, a weak solution $\bm{v}\in\mathrm{C}([0,T];\mathbf{V})\cap\mathrm{L}^{p}(0,T;\mathbf{V}_{p})$ to the problem (1.1).
Theorem 3.3.
Under the assumptions of Theorem 3.2, the solution to the problem (1.1) is unique in the class of weak solutions defined in Definition 3.1.
The proof of Theorem 3.2 is postponed to Sections 4 and 6, whereas the proof of Theorem 3.3 is carried out in Section 7.
Regarding the proof of Theorem 3.2, we proceed by distinguishing two separate cases:
$p\geq\frac{3d-4}{d}\quad\text{and}\quad\frac{2d}{d+2}<p<\frac{3d-4}{d}.$ (3.6)
This splitting is dictated by the following embedding, which refines the classical embedding (1.8) employed to establish the existence of weak solutions in [40]:
$\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,2}_{0}(\Omega)^{d})\cap\mathrm{L}^{p}(0,T;\mathrm{W}^{1,p}_{0}(\Omega)^{d})\hookrightarrow\mathrm{L}^{\rho}(0,T;\mathrm{L}^{\rho}(\Omega)^{d}),\quad\rho:=\frac{pd}{d-2},\quad d>2,$ (3.7)
which follows from the classical Sobolev embedding combined with parabolic interpolation arguments. The special cases $d=2$ and $d=p$ require less work, because the embeddings $\mathrm{W}^{1,2}_{0}(\Omega)^{d}\hookrightarrow\mathrm{L}^{\rho}(\Omega)^{d}$, when $d=2$, and $\mathrm{W}^{1,p}_{0}(\Omega)^{d}\hookrightarrow\mathrm{L}^{\rho}(\Omega)^{d}$, for $d=p$, are both valid for any $\rho\in(1,\infty)$. Under the embedding (3.7), one obtains
$t\longmapsto\int_{\Omega}\bm{v}(t)\otimes\bm{v}(t):\nabla\bm{\varphi}(t)\,\mathrm{d}\bm{x}\ \in\ \mathrm{L}^{1}(0,T),$ (3.8)
for all $\bm{v},\bm{\varphi}\in\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,2}_{0}(\Omega)^{d})\cap\mathrm{L}^{p}(0,T;\mathrm{W}^{1,p}_{0}(\Omega)^{d})$, provided that
$\frac{2}{\rho}+\frac{1}{p}\leq 1\quad\Leftrightarrow\quad p\geq\frac{3d-4}{d}.$
When $p$ satisfies the above bound, the space of test functions naturally coincides with the solution space, which simplifies the analysis considerably. In this case, the existence of weak solutions follows by standard application of the monotone operator theory [35, 40].
4. Existence: large values of $p$
In this section, we provide the proof of Theorem 3.2 in the case of
$p\geq\frac{3d-4}{d}.$ (4.1)
4.1. Approximate solutions
We construct a solution to the problem (1.1) as a limit of suitable Galerkin approximations. Let $s$ be the smallest positive integer such that
$\mathrm{W}^{s,2}_{0}(\Omega)^{d}\hookrightarrow\mathrm{W}^{1,\infty}_{0}(\Omega)^{d},\qquad s>1+\frac{d}{2},$ (4.2)
and let us consider the space $\mathbf{V}_{s}$ associated to $\mathrm{W}^{s,2}_{0}(\Omega)^{d}$ defined in Section 2. By means of separability, there exists a basis $\big\{\mathds{\bm{\psi}}_{k}\big\}_{k\in\mathbb{N}}$ of $\mathbf{V}_{s}$, formed by the eigenfunctions of a suitable spectral problem, that is, orthogonal in $\mathrm{L}^{2}(\Omega)^{d}$ and that can be made orthonormal in $\mathrm{W}^{s,2}_{0}(\Omega)^{d}$ (see [41, Theorem A.4.11]). Given $n\in\mathbb{N}$, let us consider the $n$-dimensional space $\mathbf{X}^{n}=\mathrm{span}\{\bm{\psi}_{1},\dots,\bm{\psi}_{n}\}$. For each $n\in\mathbb{N}$, we search for approximate solutions $\bm{v}_{n}\in\mathrm{C}^{1}(0,T;\mathbf{X}^{n})$ of the form
$\bm{v}_{n}(x,t)=\sum_{k=1}^{n}c_{k}^{n}(t)\bm{\psi}_{k}(x),\quad\bm{\psi}_{k}\in\mathbf{X}^{n},$ (4.3)
where the coefficients $c_{1}^{n}(t),\dots,c_{n}^{n}(t)$ are solutions of the following $n$ ordinary differential equations, derived from (3.2),
$\begin{split}&\frac{\mathrm{d}}{\,\mathrm{dt}}\mathopen{}\mathclose{{\left[\big(\bm{v}_{n}(t),\bm{\psi}_{k}\big)+\kappa\big(\nabla\bm{v}_{n}(t),\nabla\bm{\psi}_{k}\big)}}\right]+\nu\big<|\mathbf{D}(\bm{v}_{n}(t))|^{p-2}\mathbf{D}(\bm{v}_{n}(t)),\mathbf{D}(\bm{\psi}_{k})\big>\\ &=\big(\bm{f}(t),\bm{\psi}_{k}\big)+\big(\bm{v}_{n}(t)\otimes\bm{v}_{n}(t):\nabla\bm{\psi}_{k}\big),\qquad k\in\{1,\dots,n\},\end{split}$ (4.4)
supplemented with the initial conditions
$\bm{v}_{n}(0)=\bm{v}_{0,n},\quad\mbox{in}\ \Omega,$ (4.5)
where $\bm{v}_{0,n}=P^{n}(\bm{v}_{0})$, and $P^{n}$ denotes the orthogonal projection
$P^{n}:\mathbf{V}\longrightarrow\mathbf{X}^{n}$ (4.6)
so that
$\bm{v}_{n}(0,x)=\sum_{k=1}^{n}c_{k}^{n}(0)\bm{\psi}_{k}(x),\quad c_{k}^{n}(0)=c_{k,0}^{n}:=(\bm{v}_{0},\bm{\psi}_{k}),\quad k\in\{1,\dots,n\}.$
By the uniform continuity of the orthogonal projection $P^{n}$, we may assume that
$\bm{v}_{0,n}\xrightarrow[n\to\infty]{}\bm{v}_{0},\quad\text{ in }\ \mathrm{W}^{1,2}_{0}(\Omega)^{d}\cap\mathbf{V}_{p},$ (4.7)
for $p$ satisfying the condition given in (4.1). Observe that the system (4.4)-(4.5) can be written in the matrix form as follows,
$\mathbf{C}\mathbf{c}^{\prime}(t)=\mathbf{b}(t),\quad\mathbf{c}(0)=\mathbf{c}_{n,0},$ (4.8)
where $\mathbf{C}=\mathopen{}\mathclose{{\left\{a^{n}_{\ell m}}}\right\}_{\ell,m}^{n}$, $\mathbf{b}(t)=\{b_{\ell}^{n}(t)\}_{\ell=1}^{n}$, $\mathbf{c}(t)=\{c_{m}^{n}(t)\}_{m=1}^{n}$, with
$a_{\ell m}^{n}$ $:=\mathopen{}\mathclose{{\left(\bm{\psi}_{\ell},\bm{\psi}_{m}}}\right)+\kappa\mathopen{}\mathclose{{\left(\nabla\bm{\psi}_{\ell},\nabla\bm{\psi}_{m}}}\right),$
$b_{\ell}^{n}(t)$ $:=\big(\bm{f}(t),\bm{\psi}_{\ell}\big)+\big(\bm{v}_{n}(t)\otimes\bm{v}_{n}(t):\nabla\bm{\psi}_{\ell}\big)-\nu\big<|\mathbf{D}(\bm{v}_{n}(t))|^{p-2}\mathbf{D}(\bm{v}_{n}(t)),\mathbf{D}(\bm{\psi}_{\ell})\big$>,
$\mathbf{c}(t)$ $:=(c_{1}^{n}(t),\dots,c_{n}^{n}(t))\qquad\mbox{and}\qquad\mathbf{c}_{0}={(c_{1,0}^{n},\dots,c_{n,0}^{n})}.$
Taking into account that the family $\big\{\bm{\psi}_{k}\big\}_{k\in\mathbb{N}}$ is linearly independent in $\mathbf{V}$, $\big(\mathbf{C}\bm{\xi},\bm{\xi}\big)>0$, for all $\bm{\xi}\in\mathbb{R}^{n}\setminus\{0\}$, we can write (4.8) in the form
$\mathbf{c}^{\prime}(t)=\overline{\mathbf{b}(t)},\quad\mathbf{c}(0)=\mathbf{c}_{n,0},$ (4.9)
where $\overline{\mathbf{b}(t)}=\mathbf{C}^{-1}\mathbf{b}(t)$. For a globally Lipschitz-continuous coefficient function $\mathbf{b}(\cdot)$, we can use the Carathéodory theorem (see [41, Theorem A.3.4]) to prove the existence of a solution $\mathbf{c}(t)$, in a short interval $0,T^{\ast}]\subset[0,T]$, to the system (4.9).
4.2. Uniform estimates
We establish the uniform energy estimates, with respect to $n$, for the solution of the finite-dimensional approximated system corresponding to the system (1.1).
Lemma 4.1.
Assume $\bm{f}\in\mathrm{L}^{p^{\prime}}(0,T;\mathrm{L}^{p^{\prime}}(\Omega)^{d})$.
  1. (1)
    If $\bm{v}_{0}\in\mathbf{V}$ and $p>1$, then there exists an independent of $n$ positive constant $C_{1}$ such that the following estimate holds true:
    $\begin{split}&\sup_{t\in[0,T]}\mathopen{}\mathclose{{\left\{\|\bm{v}_{n}(t)\|^{2}_{2}+\kappa\|\nabla\bm{v}_{n}(t)\|_{2}^{2}}}\right\}+\nu\int_{0}^{T}\|\nabla\bm{v}_{n}(t)\|^{p}_{p}\,\mathrm{dt}\\ &\leq C_{1}\mathopen{}\mathclose{{\left(\int_{0}^{T}\|\bm{f}(t)\|^{p^{\prime}}_{p^{\prime}}\,\mathrm{dt}+\|\bm{v}_{0}\|^{2}_{2}+\|\nabla\bm{v}_{0}\|_{2}^{2}}}\right);\end{split}$ (4.10)
  2. (2)
    If $\bm{v}_{0}\in\mathbf{V}_{p}$, $2\leq d\leq 3$ and $p>1$, or $d>3$ and $p\geq d$, are verified, then there exists an independent of $n$ positive constant $C_{2}$ such that the following estimate holds true :
    $\begin{split}&\sup_{t\in[0,T]}\|\nabla\bm{v}_{n}(t)\|^{p}_{p}+\int_{0}^{T}\mathopen{}\mathclose{{\left(\|{\partial_{t}\bm{v}_{n}}(t)\|_{2}^{2}+\kappa\|\nabla{\partial_{t}\bm{v}_{n}}(t)\|^{2}_{2}}}\right)\,\mathrm{dt}\\ &\leq C_{2}\mathopen{}\mathclose{{\left(\int_{0}^{T}\|\bm{f}(t)\|^{p^{\prime}}_{p^{\prime}}\,\mathrm{dt}+\|\nabla\bm{v}_{0}\|^{p}_{p}}}\right).\end{split}$ (4.11)
Proof.
To establish (4.10) and (4.11), we multiply equation (4.4) by $c_{k}^{n}(t)$ and $\frac{\mathrm{d}c_{k}^{n}(t)}{\mathrm{d}t}$, respectively, and sum the resulting expressions over $k=1$ to $k=n$. The desired estimates then follow from suitable applications of Hölder’s, Korn’s, Sobolev’s, Young’s, and Gronwall’s inequalities. The structure of the proof parallels the approach used in [7, Lemma 3.2], though our momentum equation differs from the one considered there. We just draw the reader’s attention to statement (2), where the distinction between the cases
$2\leq d\leq 3\quad\text{with}\quad p>1\qquad\text{and}\qquad d>3\quad\text{with}\quad p\geq d$
arises from the need to control the boundedness of the convective integral term
$\int_{\Omega}\big(\bm{v}_{n}(t)\cdot\nabla\big)\bm{v}_{n}(t)\cdot\partial_{t}\bm{v}_{n}(t)\,\mathrm{d}x.$ (4.12)
When $d=2$ or $d=3$, the Ladyzhenskaya inequalities (cf. [35, Ch. 1, Lemmas 1–2]) can be applied as follows:
$\begin{split}\mathopen{}\mathclose{{\left|\int_{\Omega}\big(\bm{v}_{n}(t)\cdot\nabla\big)\partial_{t}\bm{v}_{n}(t)\cdot\bm{v}_{n}(t)\,\mathrm{d}x}}\right|&\leq\|\nabla\partial_{t}\bm{v}_{n}(t)\|_{2}\,\|\bm{v}_{n}(t)\|_{4}^{2}\\ &\leq\begin{cases}\sqrt{2}\|\nabla\partial_{t}\bm{v}_{n}(t)\|_{2}\|\bm{v}_{n}(t)\|_{2}\|\nabla\bm{v}_{n}(t)\|_{2},&d=2,\\[4.0pt] 2\|\nabla\partial_{t}\bm{v}_{n}(t)\|_{2}\|\bm{v}_{n}(t)\|_{2}^{\frac{1}{2}}\|\nabla\bm{v}_{n}(t)\|_{2}^{\frac{3}{2}},&d=3.\end{cases}\end{split}$
For $d>3$, we use Hölder’s inequality:
$\mathopen{}\mathclose{{\left|\int_{\Omega}\big(\bm{v}_{n}(t)\cdot\nabla\big)\bm{v}_{n}(t)\cdot\partial_{t}\bm{v}_{n}(t)\,\mathrm{d}x}}\right|\leq\|\bm{v}_{n}(t)\|_{2}\,\|\nabla\bm{v}_{n}(t)\|_{p}\,\|\partial_{t}\bm{v}_{n}(t)\|_{\frac{2d}{d-2}},$
which holds provided that
$\frac{1}{2}+\frac{1}{p}+\frac{1}{\frac{2d}{d-2}}\leq 1\quad\Leftrightarrow\quad p\geq d.$
Remark 4.2.
Throughout this work, we focus on the parameter range (3.5), which constitutes the central contribution of this study. The restriction on the exponent $p$ appearing in statement (2) of Lemma 4.1, specifically the case $p\geq d$ when $d>3$, is of a technical nature and is mainly related to the boundedness of the convective term (4.12). For the purposes of the existence analysis, however, it is sufficient to obtain uniform bounds on the Galerkin approximated time derivative $\partial_{t}\bm{v}_{n}$ in $\mathrm{W}^{-s,2}(\Omega)^{d}$ in order to invoke the Aubin-Dubinskiĭ compactness lemma. Therefore, the precise restriction $p\geq d$ when $d>3$ is not essential for establishing the existence of weak solutions, it mainly facilitates the technical estimates for the nonlinear term. See Remarks 4.3 and 4.4 below.
The uniform estimate (4.10) enables us to use the Continuation Principle to extend the solution $\mathbf{c}(t)$ of the initial-value problem (4.9) to the whole interval $[0,T]$.
On the other hand, as a consequence of (3.7) and (4.10), we have
$\int_{0}^{T}\|\bm{v}_{n}(t)\otimes\bm{v}_{n}(t)\|_{\frac{\rho}{2}}^{\frac{\rho}{2}}\,\mathrm{dt}\leq C,\qquad\rho:=\frac{pd}{d-2},$ (4.13)
for some positive constant $C$ not depending on $n$. Note that
$\frac{\rho}{2}>1\Leftrightarrow p>\frac{2(d-2)}{d}\quad\Leftarrow\quad p\geq\frac{3d-4}{d}\ \text{ and }\ \frac{3d-4}{d}>\frac{2(d-2)}{d}\Leftrightarrow d>0.$
From (4.11), we have
$\int_{0}^{T}\|\partial_{t}\bm{v}_{n}(t)\|_{1,2}^{2}\,\mathrm{d}t\leq C.$ (4.14)
Using the embedding $\mathrm{W}_{0}^{1,2}(\Omega)^{d}\hookrightarrow\mathrm{L}^{q}(\Omega)^{d}$, which holds for $q=\frac{2d}{d-2}$ when $d\neq 2$, or for any finite $q\geq 1$ when $d=2$, together with the embedding
$\mathrm{W}_{0}^{1,r^{\prime}}(\Omega)^{d}\hookrightarrow\mathrm{L}^{q^{\prime}}(\Omega)^{d}\ \Leftrightarrow\ \mathrm{L}^{q}(\Omega)^{d}\hookrightarrow\mathrm{W}^{-1,r}(\Omega)^{d},$
which is valid under the condition
$\mathopen{}\mathclose{{\left(\frac{2d}{d-2}}}\right)^{\prime}=q^{\prime}\leq(r^{\prime})^{\star}\ \Leftrightarrow\ r^{\prime}\geq\frac{2d}{d+4}\quad\Leftarrow\quad r^{\prime}>1,$
when $r^{\prime}<d$, or for any $r^{\prime}\in[1,\infty)$ when $r^{\prime}=d$, or any $r^{\prime}\in[1,\infty]$ when $r^{\prime}>d$, we deduce from (4.14) that
$\int_{0}^{T}\|\partial_{t}\bm{v}_{n}(t)\|_{-1,r}^{2}\,\mathrm{d}t\leq C,\qquad 1\leq r\leq\frac{2d}{d-2},$ (4.15)
for $d\neq 2$, or for any finite $r\geq 1$ when $d=2$.
Remark 4.3.
Estimate (4.10) provides uniform bounds on $\bm{v}_{n}$ in $\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,2}(\Omega)^{d})$ and in $\mathrm{L}^{p}(0,T;\mathrm{W}^{1,p}(\Omega)^{d})$, independently of $n$. The Galerkin system (4.4) allows us to write
$(\mathrm{I}-\kappa\Delta)\partial_{t}\bm{v}_{n}=P^{n}\Big(-\mathbf{div}\,(\bm{v}_{n}\otimes\bm{v}_{n})+\nu\mathbf{div}\,(|\mathbf{D}(\bm{v}_{n})|^{p-2}\mathbf{D}(\bm{v}_{n}))+\bm{f}\Big),$ (4.16)
where $P^{n}$ is the orthogonal projection defined in (4.6). As $\kappa>0$, the elliptic operator $(\mathrm{I}-\kappa\Delta)$ is invertible, and, because $P^{n}$ acts on a finite-dimensional space, $(\mathrm{I}-\kappa\Delta)^{-1}P^{n}$ is bounded. Therefore it is sufficient to estimate the right-hand side of (4.16) in $\mathrm{W}^{-s,2}(\Omega)^{d}$ to control $\partial_{t}\bm{v}_{n}$.
The convective term can be estimated, for any $\bm{\phi}\in\mathrm{W}_{0}^{s,2}(\Omega)^{d}$, as
$\mathopen{}\mathclose{{\left|\big\langle-\mathbf{div}\,(\bm{v}_{n}\otimes\bm{v}_{n}),\bm{\phi}\big\rangle}}\right|=\mathopen{}\mathclose{{\left|\int_{\Omega}(\bm{v}_{n}\cdot\nabla)\bm{v}_{n}\cdot\bm{\phi}\,\mathrm{d}x}}\right|\leq\|\bm{v}_{n}\|_{2^{\ast}}\,\|\nabla\bm{v}_{n}\|_{p}\,\|\bm{\phi}\|_{2^{\ast}_{s}},$
where $2^{\ast}=\frac{2d}{d-2}$ (for $d>2$, or arbitrarily large if $d=2$) and $2^{\ast}_{s}=\frac{2d}{d-2s}$ for $s<\frac{d}{2}$. Let us focus on the case $d>2$, once for $d=2$ the analysis is simpler. According to (4.2), $s>1+\frac{d}{2}$, and thus we can take $2^{\ast}_{s}=\infty$. Then, Hölder’s inequality requires
$\frac{d-2}{2d}+\frac{1}{p}<1\Leftrightarrow p>\frac{2d}{d+2}.$
The viscous term $\nu\,\mathbf{div}\,(|\mathbf{D}(\bm{v}_{n})|^{p-2}\mathbf{D}(\bm{v}_{n}))$ is naturally controlled in $\mathrm{W}^{-1,p^{\prime}}(\Omega)^{d}$ and the same is true for the forcing term $\bm{f}$ in view of assumption (3.4).
Hence, a simultaneous control of all terms is achieved by taking $s=p$. As a result, for any dimension $d\geq 2$, it holds that
$\partial_{t}\bm{v}_{n}\in\mathrm{L}^{p^{\prime}}(0,T;\mathrm{W}^{-s,2}(\Omega)^{d}).$ (4.17)
4.3. Passing to the limit as $n\to\infty$
In view of (4.10) and (4.11), the Banach–Alaoglu theorem guarantees the existence of functions $\bm{v}$ and $\bm{\mathcal{A}}$ such that, for some subsequences (still denoted by the same indices), we have:
$\bm{v}_{n}$ $\xrightharpoonup[n\to\infty]{\ast}\bm{v},\quad$ $\text{in}\quad\mathrm{L}^{\infty}(0,T;\mathbf{V}),$ (4.18)
$\bm{v}_{n}$ $\xrightharpoonup[n\to\infty]{\ast}\bm{v},\quad$ $\text{in}\quad\mathrm{L}^{\infty}(0,T;\mathbf{V}_{p}),$ (4.19)
$\bm{v}_{n}$ $\xrightharpoonup[n\to\infty]{}\bm{v},\quad$ $\text{in}\quad\mathrm{L}^{p}(0,T;\mathbf{V}_{p}),$ (4.20)
$\mathbf{A}(\bm{v}_{n})$ $\xrightharpoonup[n\to\infty]{}\bm{\mathcal{A}},\quad$ $\text{in}\quad\mathrm{L}^{p^{\prime}}(0,T;\mathrm{L}^{p^{\prime}}(\Omega)^{d\times d}),$ (4.21)
$\partial_{t}\bm{v}_{n}$ $\xrightharpoonup[n\to\infty]{}\partial_{t}\bm{v},\quad$ $\text{in}\quad\mathrm{L}^{2}(0,T;\mathbf{V}).$ (4.22)
Assuming $d>2$ (since the case $d=2$ is simpler), and combining (4.10) and (4.15) with the compact and continuous embeddings
$\mathrm{W}^{1,p}_{0}(\Omega)^{d}\hookrightarrow\hookrightarrow\mathrm{L}^{2}(\Omega)^{d}\hookrightarrow\mathrm{W}^{-1,r}(\Omega)^{d},\quad p>\frac{2d}{d+2},\quad r\leq\frac{2d}{d-2},\quad d\neq 2,$ (4.23)
we may apply Lemma 2.2-(2) with $s=0$, $p=2$, $p_{0}=2$, and $p_{1}=p$. This yields the existence of a function $\bm{v}$ such that, for some subsequence (again not relabeled),
$\bm{v}_{n}\xrightarrow[n\to\infty]{}\bm{v},\quad\text{in}\quad\mathrm{L}^{2}(Q_{T})^{d}.$ (4.24)
By the Riesz–Fischer theorem, there exists a subsequence, still denoted by the same index, such that
$\bm{v}_{n}\xrightarrow[n\to\infty]{}\bm{v},\quad\text{a.e. in }Q_{T},$
and consequently,
$\bm{v}_{n}\otimes\bm{v}_{n}\xrightarrow[n\to\infty]{}\bm{v}\otimes\bm{v},\quad\text{a.e. in }Q_{T}.$ (4.25)
Now, in view of (4.13) and (4.25), a classical result (see e.g. [40, Lemma 1.3]) yields
$\bm{v}_{n}\otimes\bm{v}_{n}\xrightharpoonup[n\to\infty]{}\bm{v}\otimes\bm{v},\quad\text{in}\quad\mathrm{L}^{\frac{\rho}{2}}(0,T;\mathrm{L}^{\frac{\rho}{2}}(\Omega)^{d\times d}).$ (4.26)
Remark 4.4.
We can avoid the use of estimate (4.11), which requires the restriction $p\geq d$ when $d>3$ and is invoked in (4.14) for the application of the Aubin–Dubinskiĭ lemma (Lemma 2.2) to obtain the strong convergence (4.24). Instead, we can follow the argument outlined in Remark 4.3. Indeed, the uniform bounds provided by (4.10) and (4.17) are sufficient to apply the Aubin–Dubinskiĭ lemma directly, thereby guaranteeing the strong convergence of the Galerkin approximations.
4.4. Application of the Minty trick
Multiplying (4.4) by $c_{k}^{n}\in\mathrm{C}^{1}(0,T)$, with $k\in\{1,\dots,n\}$, and adding up the resulting equation from $k=1$ to $n$, and integrating over $(0,T)$, one has
$\begin{split}&\!\int_{0}^{T}\!\!\!\!\int_{\Omega}\big({\partial_{t}\bm{v}_{n}}\cdot\bm{\psi}_{j}+\kappa\nabla{\partial_{t}\bm{v}_{n}}:\nabla\bm{\psi}_{j}\big)\,\mathrm{d}\bm{x}\mathrm{d}t+\nu\!\int_{0}^{T}\!\!\!\!\int_{\Omega}|\mathbf{D}(\bm{v}_{n})|^{p-2}\mathbf{D}(\bm{v}_{n}):\mathbf{D}(\bm{\psi}_{j})\,\mathrm{d}\bm{x}\mathrm{d}t\\ &=\!\int_{0}^{T}\!\!\!\!\int_{\Omega}\bm{f}\cdot\bm{\psi}_{j}\,\mathrm{d}\bm{x}\mathrm{d}t+\!\int_{0}^{T}\!\!\!\!\int_{\Omega}\bm{v}_{n}\otimes\bm{v}_{n}:\nabla\bm{\psi}_{j}\,\mathrm{d}\bm{x}\mathrm{d}t,\end{split}$ (4.27)
for any function
$\bm{\psi}_{j}(\bm{x},t)=\sum\limits_{k=1}^{j}{c_{k}^{j}(t)}\bm{\psi}_{k}(x),\qquad\bm{\psi}_{k}\in\mathbf{X}^{j},\quad k\in\{1,\dots,j\},\quad j\in\{1,\dots,n\}.$
In particular, taking $\bm{\psi}_{j}=\bm{v}_{n}$ in (4.27), we have
$\begin{split}&\|\bm{v}_{n}(T)\|^{2}_{2}+\kappa\|\nabla\bm{v}_{n}(T)\|_{2}^{2}+2\nu\!\int_{0}^{T}\!\!\!\!\int_{\Omega}|\mathbf{D}(\bm{v}_{n})|^{p}\,\mathrm{d}\bm{x}\mathrm{d}t\\ &=\|\bm{v}_{0,n}\|^{2}_{2}+\kappa\|\nabla\bm{v}_{0,n}\|_{2}^{2}+2\!\int_{0}^{T}\!\!\!\!\int_{\Omega}\bm{f}\cdot\bm{v}_{n}\,\mathrm{d}\bm{x}\mathrm{d}t.\end{split}$ (4.28)
On the other hand, using the convergence results (4.18)-(4.22) and (4.26), we can pass to the limit as $n\to\infty$ in (4.27) so that
$\begin{split}&\!\int_{0}^{T}\!\!\!\!\int_{\Omega}\big({\partial_{t}\bm{v}}\cdot\bm{\psi}_{j}+\kappa\nabla{\partial_{t}\bm{v}}:\nabla\bm{\psi}_{j}\big)\,\mathrm{d}\bm{x}\mathrm{d}t+\nu\!\int_{0}^{T}\!\!\!\!\int_{\Omega}\bm{\mathcal{A}}:\nabla\bm{\psi}_{j}\,\mathrm{d}\bm{x}\mathrm{d}t\\ &=\!\int_{0}^{T}\!\!\!\!\int_{\Omega}\bm{f}\cdot\bm{\psi}_{j}\,\mathrm{d}\bm{x}\mathrm{d}t+\!\int_{0}^{T}\!\!\!\!\int_{\Omega}\bm{v}\otimes\bm{v}:\nabla\bm{\psi}_{j}\,\mathrm{d}\bm{x}\mathrm{d}t,\end{split}$ (4.29)
for any such function $\bm{\psi}_{j}$. Since the functions $\mathopen{}\mathclose{{\left\{\bm{\psi}_{j}}}\right\}_{j\in\mathbb{N}}$ are dense in $\mathrm{L}^{\infty}(0,T;\mathbf{V})\cap\mathrm{L}^{p}(0,T;\mathbf{V}_{p})$, we can put $\bm{\psi}_{j}=\bm{v}$ in (4.29), which yields
$\begin{split}&\|\bm{v}(T)\|^{2}_{2}+\kappa\|\nabla\bm{v}(T)\|_{2}^{2}+2\nu\!\int_{0}^{T}\!\!\!\!\int_{\Omega}\bm{\mathcal{A}}:\nabla\bm{v}\,\mathrm{d}\bm{x}\mathrm{d}t\\ &=\|\bm{v}_{0}\|^{2}_{2}+\kappa\|\nabla\bm{v}_{0}\|_{2}^{2}+2\!\int_{0}^{T}\!\!\!\!\int_{\Omega}\bm{f}\cdot\bm{v}\,\mathrm{d}\bm{x}\mathrm{d}t.\end{split}$ (4.30)
Now, we subtract (4.30) from (4.28), and thus we have
$\begin{split}&\!\int_{0}^{T}\!\!\!\!\int_{\Omega}|\mathbf{D}(\bm{v}_{n})|^{p}\,\mathrm{d}\bm{x}\mathrm{d}t-\!\int_{0}^{T}\!\!\!\!\int_{\Omega}\bm{\mathcal{A}}:\mathbf{D}(\bm{v})\,\mathrm{d}\bm{x}\mathrm{d}t=:\frac{1}{2\nu}\big(I_{1}+I_{2}+I_{3}\big),\end{split}$ (4.31)
where
$I_{1}:=\|\bm{v}(T)\|^{2}_{2}+\kappa\|\nabla\bm{v}(T)\|_{2}^{2}-\big(\|\bm{v}_{n}(T)\|^{2}_{2}+\kappa\|\nabla\bm{v}_{n}(T)\|_{2}^{2}\big),$
$I_{2}:=\|\bm{v}_{0,n}\|^{2}_{2}+\kappa\|\nabla\bm{v}_{0,n}\|_{2}^{2}-\big(\|\bm{v}_{0}\|^{2}_{2}+\kappa\|\nabla\bm{v}_{0}\|_{2}^{2}\big),$
$I_{3}:=2\!\int_{0}^{T}\!\!\!\!\int_{\Omega}\bm{f}\cdot\big(\bm{v}_{n}-\bm{v}\big)\,\mathrm{d}\bm{x}\mathrm{d}t.$
By the monotonicity of the operator $\mathbf{A}(\bm{v})=|\mathbf{D}(\bm{v})|^{p-2}\mathbf{D}(\bm{v})$ (see (2.3)), there holds
$\begin{split}\!\int_{0}^{T}\!\!\!\!\int_{\Omega}|\mathbf{D}(\bm{v}_{n})|^{p}\,\mathrm{d}\bm{x}\mathrm{d}t&\geq\!\int_{0}^{T}\!\!\!\!\int_{\Omega}|\mathbf{D}(\bm{\psi})|^{p-2}\mathbf{D}(\bm{\psi}):\mathbf{D}(\bm{v}_{n}-\bm{\psi})\,\mathrm{d}\bm{x}\mathrm{d}t\\ &\quad+\!\int_{0}^{T}\!\!\!\!\int_{\Omega}|\mathbf{D}(\bm{v}_{n})|^{p-2}\mathbf{D}(\bm{v}_{n}):\mathbf{D}(\bm{\psi})\,\mathrm{d}\bm{x}\mathrm{d}t,\end{split}$ (4.32)
for any function $\bm{\psi}\in\overline{\mathrm{span}\{\bm{\psi}_{1},\dots,\bm{\psi}_{n}\}}$. Then, combining (4.31) with (4.32), we have
$\begin{split}&\!\int_{0}^{T}\!\!\!\!\int_{\Omega}|\mathbf{D}(\bm{\psi})|^{p-2}\mathbf{D}(\bm{\psi}):\mathbf{D}(\bm{v}_{n}-\bm{\psi})\,\mathrm{d}\bm{x}\mathrm{d}t+\!\int_{0}^{T}\!\!\!\!\int_{\Omega}|\mathbf{D}(\bm{v}_{n})|^{p-2}\mathbf{D}(\bm{v}_{n}):\mathbf{D}(\bm{\psi})\,\mathrm{d}\bm{x}\mathrm{d}t\\ &\qquad-\!\int_{0}^{T}\!\!\!\!\int_{\Omega}\bm{\mathcal{A}}:\mathbf{D}(\bm{v})\,\mathrm{d}\bm{x}\mathrm{d}t\\ &\leq\frac{1}{2\nu}\mathopen{}\mathclose{{\left(I_{1}+I_{2}+I_{3}}}\right).\end{split}$ (4.33)
Using the convergence results (4.20) and (4.21), we can pass (4.33) to the limit, as $n\to\infty$, so that
$\begin{split}&\!\int_{0}^{T}\!\!\!\!\int_{\Omega}\big(|\mathbf{D}(\bm{\psi})|^{p-2}\mathbf{D}(\bm{\psi})-\bm{\mathcal{A}}\big):\mathbf{D}(\bm{v}-\bm{\psi})\leq C\lim_{n\to\infty}\mathopen{}\mathclose{{\left(I_{1}+I_{2}+I_{3}}}\right),\end{split}$ (4.34)
for some positive constant $C=C(\nu)$. By (4.18), the properties of $\limsup$ and $\liminf$, and the lower semicontinuity of the norm, one has
$\lim\limits_{n\to\infty}I_{1}\leq\limsup\limits_{n\to\infty}I_{1}\leq{\|\bm{v}(T)\|^{2}_{2}+\kappa\|\nabla\bm{v}(T)\|_{2}^{2}}-\liminf\limits_{n\to\infty}\big(\|\bm{v}_{n}(T)\|^{2}_{2}+\kappa\|\nabla\bm{v}_{n}(T)\|_{2}^{2}\big)\leq 0.$
By construction (see (4.5)-(4.7)),
$\lim\limits_{n\to\infty}I_{2}=0.$
Due to assumption (3.4) and (4.20),
$\lim\limits_{n\to\infty}I_{3}=0.$
The above convergences allow us to write (4.34) as follows,
$\begin{split}&\!\int_{0}^{T}\!\!\!\!\int_{\Omega}\big(|\mathbf{D}(\bm{\psi})|^{p-2}\mathbf{D}(\bm{\psi})-\bm{\mathcal{A}}\big):\mathbf{D}(\bm{v}-\bm{\psi})\,\mathrm{d}\bm{x}\mathrm{d}t\leq 0.\end{split}$ (4.35)
Taking $\bm{\psi}=\bm{v}\pm\delta\bm{u}$, with $\delta>0$ and $\bm{u}\in\mathrm{L}^{\infty}(0,T;\mathbf{V})\cap\mathrm{L}^{p}(0,T;\mathbf{V}_{p})$, in (4.35), we obtain
$\begin{split}&\mp\delta\!\int_{0}^{T}\!\!\!\!\int_{\Omega}\big(|\mathbf{D}(\bm{v}\pm\delta\bm{u})|^{p-2}\mathbf{D}(\bm{v}\pm\delta\bm{u})-\bm{\mathcal{A}}\big):\mathbf{D}(\bm{u})\,\mathrm{d}\bm{x}\mathrm{d}t\leq 0.\end{split}$
Letting $\delta\longrightarrow 0$, we see that it must be
$\mp\!\int_{0}^{T}\!\!\!\!\int_{\Omega}\big(|\mathbf{D}(\bm{v})|^{p-2}\mathbf{D}(\bm{v})-\bm{\mathcal{A}}\big):\mathbf{D}(\bm{u})\,\mathrm{d}\bm{x}\mathrm{d}t=0,$
for all $\bm{u}\in\mathrm{L}^{\infty}(0,T;\mathbf{V})\cap\mathrm{L}^{p}(0,T;\mathbf{V}_{p})$. Hence
$\bm{\mathcal{A}}=|\mathbf{D}(\bm{v})|^{p-2}\mathbf{D}(\bm{v}),\quad\mbox{a.e. in}\ \ Q_{T}.$ (4.36)
4.5. Conclusion
We can see that (1)-(2) of Definition 3.1 are easily verified due to (4.18), (4.20) and (4.22). On the other hand, multiplying (4.4) by $\psi\in\mathrm{C}^{\infty}(0,T)$, with $\psi(T)=0$, integrating over $(0,T)$, using integration by parts in time, using (4.5), and the convergence results (4.7), (4.18), (4.21), (4.26) (observing here (3.8) for the choice of $\rho$ in (3.7)) and (4.36), we obtain (3.2) for any $\bm{\varphi}=\psi\bm{\psi}_{k}$, with $\bm{\psi}_{k}\in\mathbf{X}^{n}$, $\psi\in\mathrm{C}^{\infty}(0,T)$, and $\psi(T)=0$. Thus, by continuity and density, (3.2) holds true for any $\bm{\varphi}\in\mathbf{C}^{\infty}_{0,\operatorname{div}}(Q_{T})$. Finally, since $\partial_{t}\bm{v}\in\mathrm{L}^{2}(0,T;\mathbf{V})$ and $\bm{v}\in\mathrm{L}^{2}(0,T;\mathbf{V})$, we have $\bm{v}\in\mathrm{W}^{1,2}(0,T;\mathbf{V})\subset\mathrm{C}([0,T];\mathbf{V})$. Consequently, the initial condition $\bm{v}(0)=\bm{v}_{0}$, stated in (3) of Definition 3.1, is meaningful. This concludes the proof of Theorem 3.2 under assumption (4.1).
In the rest of this work, we focus on the most challenging case in the proof of Theorem 3.2, that is,
$\frac{2d}{d+2}<p<\frac{3d-4}{d}.$ (4.37)
This intermediate range of the power-law exponent $p$ is significantly more delicate because the embedding stated in (3.7) no longer holds. As a consequence, additional analytical tools are required to handle the nonlinear terms. In particular, it becomes necessary to perform a refined decomposition of the pressure into multiple components in order to compensate for the loss of compactness and to secure the weak convergence properties essential to the existence argument. The next section is devoted to developing the technical aspects of this pressure decomposition, which will play a crucial role in the construction of weak solutions for the case of $p$ satisfying (4.37).
5. Revisiting the pressure decomposition technique
In this section, we look at the pressure term that appears in our model and will decompose it so that each part matches a specific term in the momentum equation. The time derivative (inertia) and Voigt terms are grouped into one part of the pressure, while the viscous and convective terms are linked to two other pressure parts that are similar but have different roles. The key idea behind the following result was originally introduced in [51] and later used in [21], building on the original insight.
Theorem 5.1.
Consider $\bm{v}\in\mathrm{C}_{\rm{w}}([0,T];\mathbf{V})$, and, for $m_{1},\,m_{2}\in(1,\infty)$, let
[…]
Assume further that condition (3.3) holds. In addition, suppose the following identity is satisfied,
[…]
for all $\bm{\varphi}\in\mathbf{C}^{\infty}_{0}(Q_{T})$ with $\operatorname{div}\bm{\varphi}=0$.
Then, there exist $\pi_{1}\in\mathrm{L}^{m_{1}}(0,T;\mathrm{L}^{m_{1}}(\Omega))$ and $\pi_{2}\in\mathrm{L}^{m_{2}}(0,T;\mathrm{W}^{1,m_{2}}(\Omega))$, and there exists $\pi_{h}\in\mathrm{C}_{\rm{w}}([0,T];\mathrm{W}^{1,2}(\Omega))$ such that
[…]
for all $\bm{\varphi}\in\mathbf{C}^{\infty}(Q_{T})$ with $\operatorname{supp}(\bm{\varphi})\Subset\Omega\times[0,T)$.
In addition, we have
$-\Delta\pi_{h}$ $=0,\quad\pi_{h}(0)=0,$ (5.3)
$\|\pi_{h}(t)\|_{1,2}$ $\leq C_{h}\big(\|\bm{v}(t)-\bm{v}(0)\|_{2}+\kappa\|\nabla\bm{v}(t)-\nabla\bm{v}(0)\|_{2}\big),$ (5.4)
[…]
for some positive constants $C_{h}=C(d,\Omega)$ and $C_{i}=C(m_{i},d,\Omega)$ for $i\in\{1,2,3\}$.
It should be noted that if we combine all the components of the decomposed pressure, we obtain
$\pi(t)=\pi_{h}(t)+\int_{0}^{t}\pi_{1}(s)\,\mathrm{d}s+\int_{0}^{t}\pi_{2}(s)\,\mathrm{d}s,$
which implies that $\pi\in\mathrm{L}^{\infty}(0,T;\mathrm{L}^{\gamma}(\Omega))$.
Proof of Theorem 5.1.
The proof has been divided into several steps.
Step 1: Let us define a bilinear form $\mathcal{Z}$ on the domain $\mathcal{D}(\mathcal{Z}):=\mathrm{W}_{0}^{1,2}(\Omega)^{d}\subset\mathrm{L}^{2}(\Omega)^{d}$ by
$\mathcal{Z}(\bm{v},\bm{u}):=\int_{\Omega}\bm{v}\cdot\bm{u}\,\mathrm{d}\bm{x}+\kappa\int_{\Omega}\nabla\bm{v}:\nabla\bm{u}\,\mathrm{d}\bm{x},$
where $\kappa>0$ is the relaxation time constant in the Navier-Stokes-Voigt model.
We begin by observing that $\mathcal{Z}$ is well-defined and bilinear on $\mathrm{W}_{0}^{1,2}(\Omega)^{d}$, since both the $\mathrm{L}^{2}$-inner product and the (Frobenius) inner product of gradients are well-defined for functions in $\mathrm{W}_{0}^{1,2}(\Omega)^{d}$.
Next, we consider the norm induced by this bilinear form, given by
$\mathcal{Z}(\bm{v},\bm{v})^{\frac{1}{2}}=\mathopen{}\mathclose{{\left(\|\bm{v}\|_{2}^{2}+\kappa\|\nabla\bm{v}\|_{2}^{2}}}\right)^{\frac{1}{2}}.$
This defines a norm on $\mathrm{W}_{0}^{1,2}(\Omega)^{d}$ equivalent to the standard $\mathrm{W}^{1,2}$-norm, due to the Poincaré inequality (which applies because of the zero boundary conditions). Hence, $\mathrm{W}_{0}^{1,2}(\Omega)^{d}$ is complete with respect to the $\mathcal{Z}$-norm. This completeness implies that the bilinear form $\mathcal{Z}$ is closed. Moreover, it follows immediately from its definition that $\mathcal{Z}$ is a symmetric and positive form.
Given that $\mathcal{Z}$ is a densely defined, closed, symmetric, and positive bilinear form on the Hilbert space $\mathrm{L}^{2}(\Omega)^{d}$, we can apply [49, Lemma II.3.2.1] to conclude the existence of a unique positive self-adjoint operator $\mathrm{A}:\mathrm{D}(\mathrm{A})\to\mathrm{L}^{2}(\Omega)^{d}$, with dense domain $\mathrm{D}(\mathrm{A})\subset\mathrm{W}_{0}^{1,2}(\Omega)^{d}$, such that
$\int_{\Omega}\mathrm{A}(\bm{v})\cdot\bm{u}\,\mathrm{d}\bm{x}=\mathcal{Z}(\bm{v},\bm{u})=\int_{\Omega}\bm{v}\cdot\bm{u}\,\mathrm{d}\bm{x}+\kappa\int_{\Omega}\nabla\bm{v}:\nabla\bm{u}\,\mathrm{d}\bm{x},$
for all $\bm{v}\in\mathrm{D}(\mathrm{A}),\ \bm{u}\in\mathrm{W}_{0}^{1,2}(\Omega)^{d}$.
Fixing an arbitrary test function $\bm{u}\in\mathrm{C}_{0}^{\infty}(\Omega)^{d}$, we observe that the relation
$\mathrm{A}(\bm{v})=(\mathrm{I}-\kappa\Delta)(\bm{v})=(\mathrm{I}-\kappa\mathbf{div}\,\nabla)(\bm{v})$
holds in the sense of distributions. Based on this (formal) identity, we define the operator $\mathrm{A}:=\mathrm{I}-\kappa\Delta$. Thus, the operator $\mathrm{A}=\mathrm{I}-\kappa\Delta:\mathrm{D}(\mathrm{I}-\kappa\Delta)\to\mathrm{L}^{2}(\Omega)^{d}$ is characterized by the relation
$\int_{\Omega}(\mathrm{I}-\kappa\Delta)(\bm{v})\cdot\bm{u}\,\mathrm{d}\bm{x}=\mathcal{Z}(\bm{v},\bm{u})=\int_{\Omega}\bm{v}\cdot\bm{u}\,\mathrm{d}\bm{x}+\kappa\int_{\Omega}\nabla\bm{v}:\nabla\bm{u}\,\mathrm{d}\bm{x},$ (5.8)
for all $\bm{v}\in\mathrm{D}(\mathrm{I}-\kappa\Delta)$ and $\bm{u}\in\mathrm{W}_{0}^{1,2}(\Omega)^{d}$. The domain of $\mathrm{A}$ is given by the set
$\mathrm{D}(\mathrm{I}-\kappa\Delta)=\mathopen{}\mathclose{{\left\{\bm{v}\in\mathrm{W}_{0}^{1,2}(\Omega)^{d}:\bm{u}\mapsto\mathcal{Z}(\bm{v},\bm{u})\text{ is continuous with respect to the }\mathrm{L}^{2}\text{-norm}}}\right\}.$
That is, $\bm{v}\in\mathrm{D}(\mathrm{I}-\kappa\Delta)$ if and only if the linear functional $\mathcal{Z}(\bm{v},\cdot)$ extends continuously to all of $\mathrm{L}^{2}(\Omega)^{d}$, meaning there exists a function $\mathrm{A}(\bm{v})\in\mathrm{L}^{2}(\Omega)^{d}$ such that the identity (5.8) holds for all $\bm{u}\in\mathrm{W}_{0}^{1,2}(\Omega)^{d}$.
We can now reformulate equation (5.1) into a weak form that is well-suited for the pressure decomposition we aim to develop. Specifically, making use of identity (5.8), we consider (5.1) in the following form:
[…]
for all $\bm{\varphi}\in\mathbf{C}^{\infty}_{0}(Q_{T})$ with $\mathbf{div}\,\bm{\varphi}=0$. This formulation sets the stage for decomposing the pressure into distinct components associated with different terms in the momentum equation.
Using (5.9) in conjunction with Lemma 2.3, and proceeding as in the proof of [51, Theorem 2.6], we obtain the existence of a (unique) pressure $\pi\in\mathrm{C}_{\mathrm{w}}([0,T];\mathrm{L}^{1}(\Omega))$, with
$\int_{\Omega}\pi(t)\mathrm{d}x=0,\quad\text{ for all }\ t\in[0,T],$ (5.10)
and such that
$\int_{\Omega}(\mathrm{I}-\kappa\Delta)\big(\bm{v}(t)-\bm{v}(0)\big)\cdot\bm{\psi}\mathrm{d}\bm{x}+\int_{\Omega}\big(\widetilde{\mathbfcal{G}}_{1}(t)+\widetilde{\mathbfcal{G}}_{2}(t)\big):\nabla\bm{\psi}\mathrm{d}\bm{x}=\int_{\Omega}\pi(t)\operatorname{div}\bm{\psi}\mathrm{d}\bm{x},$ (5.11)
for all $\bm{\psi}\in\mathrm{W}_{0}^{1,2}(\Omega)^{d}$ and for all $t\in[0,T]$, where
[…]
Step 2: Next, for $i\in\{1,2\}$ and $t\in[0,T]$, let $\bm{u}_{i}(t)\in\mathbf{V}_{m_{i}}$ be the unique solution (cf. [25, Theorem IV.6.1]) to the following Stokes problem,
$\mathopen{}\mathclose{{\left\{\begin{aligned} -\Delta\bm{u}_{i}(t)+\nabla\widetilde{\pi}_{i}(t)&=\mathbf{div}\,\widetilde{\mathbfcal{G}}_{i}(t),\qquad&&\text{in }\Omega,\\ \mathbf{div}\,\bm{u}_{i}(t)&=0,\qquad&&\text{in }\Omega,\\ \bm{u}_{i}(t)&=0,\qquad&&\text{on }\partial\Omega.\end{aligned}}}\right$. (5.13)
In view of Lemma 2.3, for each $i\in\{1,2\}$,
$\widetilde{\pi}_{i}(t)\in\mathrm{L}^{m_{i}}(\Omega),\text{ and }\int_{\Omega}\widetilde{\pi}_{i}(t)\,\mathrm{d}x=0,\quad\text{ for all }\ t\in[0,T].$ (5.14)
Proceeding as in the proof of [21, Theorem 2.2], we show that, in fact, for each $i\in\{1,2\}$,
$\widetilde{\pi}_{i}\in\mathrm{L}^{m_{i}}(0,T;\mathrm{L}^{m_{i}}(\Omega)),$
and
[…]
for some positive constants $C_{i}$, $i\in\{1,2\}$. Setting
$\pi_{i}:=\partial_{t}\widetilde{\pi}_{i},\quad i\in\{1,2\},$ (5.15)
we immediately obtain (5.5)-(5.6). The proof of (5.7) is considerably longer, so we refer the reader to the proof of [51, Lemma 2.3], where it is presented in full detail.
Now, for $t\in[0,T]$, let $\bm{u}_{h}(t)\in\mathbf{V}\cap\mathrm{W}^{2,2}(\Omega)^{d}$ be the unique solution (cf. [25, Theorem IV.6.1]) to the following Stokes problem,
$\mathopen{}\mathclose{{\left\{\begin{aligned} -\Delta\bm{u}_{h}(t)+\nabla\pi_{h}(t)&=-(\mathrm{I}-\kappa\Delta)\big(\bm{v}(t)-\bm{v}(0)\big),\qquad&&\text{in }\Omega,\\ \mathbf{div}\,\bm{u}_{h}(t)&=0,\qquad&&\text{in }\Omega,\\ \bm{u}_{h}(t)&=0,\qquad&&\text{on }\partial\Omega.\end{aligned}}}\right$. (5.16)
Moreover,
$\int_{\Omega}\pi_{h}(t)\,\mathrm{d}x$ $=0,$ (5.17)
$\|\pi_{h}(t)\|_{1,2}$ $\leq C_{h}\|(\mathrm{I}-\kappa\Delta)(\bm{v}(t)-\bm{v}(0))\|_{2}$
$\leq C_{h}\mathopen{}\mathclose{{\left(\|\bm{v}(t)-\bm{v}(0)\|_{2}+\kappa\|\nabla\bm{v}(t)-\nabla\bm{v}(0)\|_{2}}}\right).$ (5.18)
The latter inequality implies (5.4). In particular, (5.3) holds, and we have
$\sup_{t\in[0,T]}\|\pi_{h}(t)\|_{1,2}\leq C_{h}\bigg(\sup_{t\in[0,T]}\mathopen{}\mathclose{{\left\{\|\bm{v}(t)\|_{2}^{2}+\kappa\|\nabla\bm{v}(t)\|_{2}^{2}}}\right\}+\|\bm{v}_{0}\|_{2}^{2}+\kappa\|\nabla\bm{v}_{0}\|_{2}^{2}\bigg),$ (5.19)
whenever (3.3) holds.
Then, taking
$\bm{v}(t):=\bm{v}_{1}(t)+\bm{v}_{2}(t)+\bm{v}_{h}(t),$
we obtain from (5.11),
$-(\mathrm{I}-\kappa\Delta)\big(\bm{v}(t)-\bm{v}(0)\big)+\mathbf{div}\,\big(\widetilde{\mathbfcal{G}}_{1}(t)-\widetilde{\mathbfcal{G}}_{2}(t)\big)=\nabla\pi(t).$ (5.20)
Combining this with the three previous Stokes problems, we obtain another Stokes problem, now with zero right-hand side:
$\mathopen{}\mathclose{{\left\{\begin{aligned} -\Delta\bm{v}(t)+\nabla\big(\widetilde{\pi}_{1}(t)+\widetilde{\pi}_{2}(t)+\pi_{h}(t)-\pi(t)\big)&=0,\qquad&&\text{in }\Omega,\\ \mathbf{div}\,\bm{v}(t)&=0,\qquad&&\text{in }\Omega,\\ \bm{v}(t)&=0,\qquad&&\text{on }\partial\Omega.\end{aligned}}}\right$. (5.21)
It is well-known that for this Stokes problem with zero forcing and homogeneous Dirichlet boundary conditions, one has
$\bm{v}(t)=0,\text{ and }\pi(t)=\widetilde{\pi}_{1}(t)+\widetilde{\pi}_{2}(t)+\pi_{h}(t).$
For $t\in(0,T)$, we insert $\bm{\psi}(\cdot,t)=\partial_{t}\bm{\varphi}(\cdot,t)$, with $\operatorname{supp}(\bm{\varphi})\Subset\Omega\times[0,T)$, in (5.11), and integrate the resulting equation over $(0,T)$ to obtain
$\begin{split}&\int_{0}^{T}\int_{\Omega}(\mathrm{I}-\kappa\Delta)\big(\bm{v}(t)-\bm{v}(0)\big)\cdot\partial_{t}\bm{\varphi}\,\mathrm{d}x\,\mathrm{d}t+\int_{0}^{T}\int_{\Omega}\big(\widetilde{\mathbfcal{G}}_{1}(t)+\widetilde{\mathbfcal{G}}_{2}(t)\big):\nabla\partial_{t}\bm{\varphi}\,\mathrm{d}x\,\mathrm{d}t\\ &=\int_{0}^{T}\int_{\Omega}\big(\widetilde{\pi}_{1}(t)+\widetilde{\pi}_{2}(t)+\pi_{h}(t)\big)\mathbf{div}\,\partial_{t}\bm{\varphi}\,\mathrm{d}x\,\mathrm{d}t.\end{split}$ (5.22)
Finally, integrating this equation by parts in time and invoking (5.8), (5.12), (5.14), (5.15), and (5.17), we obtain the decomposition identity (5.2). ∎
6. Existence: small values of $p$
In this section, the proof of Theorem 3.2 when the power-law exponent satisfies (4.37) is carried out. We begin by regularizing problem (1.1) so that the results established in Section 4 remain applicable for the regularized problem (6.3) below. We choose a Lebesgue exponent $\beta$ such that
$\beta\geq\max\mathopen{}\mathclose{{\left\{\frac{3d-4}{d},p}}\right\}.$ (6.1)
For technical convenience in the arguments that follow (see (6.22) below), we also impose an upper bound on $\beta$. As a result, the admissible range of $\beta$ in (6.1) is restricted to
$\max\mathopen{}\mathclose{{\left\{\frac{3d-4}{d},p}}\right\}\leq\beta\leq d.$ (6.2)
6.1. Regularized problem
For $\beta$ in the conditions of (6.2) and for each $n\in\mathbb{N}$, we consider the following regularized problem in the spirit of [53],
[…]
depending on the forcing term $\bm{f}$, and on the initial velocity $\bm{v}_{0}$ in the sense that
$\bm{v}_{0,n}\xrightarrow[n\to\infty]{}\bm{v}_{0},\quad\text{in }\ \mathbf{V}.$ (6.4)
The operator $\mathbf{A}$ is defined in (2.4), and the operator $\mathbf{B}$ is defined by
$\mathbf{B}(\bm{v}):=|\mathbf{D}(\bm{v})|^{\beta-2}\mathbf{D}(\bm{v}),$ (6.5)
for $\beta$ given in (6.2). Note that, under assumption (6.1), $\beta\geq p$ and therefore $\mathrm{L}^{\beta}(0,T;\mathbf{V}_{\beta})\subset\mathrm{L}^{p}(0,T;\mathbf{V}_{p})$, which in turn implies that any weak solution to the problem (6.3) that is in $\mathrm{L}^{\beta}(0,T;\mathbf{V}_{\beta})$ is also in $\mathrm{L}^{p}(0,T;\mathbf{V}_{p})$.
Definition 6.1.
Let the conditions of Definition 3.1 be satisfied and assume (6.1) holds. We say that $\bm{v}_{n}$ is a weak solution to the problem (6.3), if in addition to (2)-(3) of Definition 3.1:
  1. (1’)
    $\bm{v}_{n}\in\mathrm{L}^{\infty}(0,T;\mathbf{V})\cap\mathrm{L}^{\beta}(0,T;\mathbf{V}_{\beta})$;
  2. (4’)
    The following identity,
    $\begin{split}&-\!\int_{0}^{T}\!\!\!\!\int_{\Omega}\bm{v}_{n}\cdot\partial_{t}\bm{\varphi}\,\mathrm{d}\bm{x}\mathrm{d}t-\kappa\!\int_{0}^{T}\!\!\!\!\int_{\Omega}\nabla\bm{v}_{n}:\nabla\partial_{t}\bm{\varphi}\,\mathrm{d}\bm{x}\mathrm{d}t+\nu\!\int_{0}^{T}\!\!\!\!\int_{\Omega}|\mathbf{D}(\bm{v}_{n})|^{p-2}\mathbf{D}(\bm{v}_{n}):\mathbf{D}(\bm{\varphi})\,\mathrm{d}\bm{x}\mathrm{d}t\\ &\quad+\frac{1}{n}\!\int_{0}^{T}\!\!\!\!\int_{\Omega}|\mathbf{D}(\bm{v}_{n})|^{\beta-2}\mathbf{D}(\bm{v}_{n}):\mathbf{D}(\bm{\varphi})\,\mathrm{d}\bm{x}\mathrm{d}t\\ &=\int_{\Omega}\bm{v}_{0}\cdot\bm{\varphi}(0)\,\mathrm{d}\bm{x}+\kappa\int_{\Omega}\nabla\bm{v}_{0}:\nabla\bm{\varphi}(0)\,\mathrm{d}\bm{x}+\!\int_{0}^{T}\!\!\!\!\int_{\Omega}\bm{f}\cdot\bm{\varphi}\,\mathrm{d}\bm{x}\mathrm{d}t+\!\int_{0}^{T}\!\!\!\!\int_{\Omega}\bm{v}_{n}\otimes\bm{v}_{n}:\mathds{\nabla}\bm{\varphi}\,\mathrm{d}\bm{x}\mathrm{d}t,\end{split}$ (6.6)
    holds for all $\bm{\varphi}\in\mathbf{C}^{\infty}_{\operatorname{div}}(Q_{T})$ with $\operatorname{supp}(\bm{\varphi})\Subset\Omega\times[0,T)$.
Since the exponent $\beta$ satisfies (6.1), it also meets condition (4.1). Therefore, the existence of a weak solution $\bm{v}_{n}$ to the problem (6.3), in the sense of Definition 6.1, can be directly inferred from the proof given in Section 4. In this case only an adaptation of assumption (3.3) is required,
$\bm{v}_{0}\in\mathbf{V}\cap\mathbf{V}_{\beta}.$ (6.7)
Note that assumption (3.4) is still usable in the present case, because, as $\beta\geq p$, we have $\mathrm{L}^{p^{\prime}}(0,T;\mathrm{L}^{p^{\prime}}(\Omega)^{d})\subset\mathrm{L}^{\beta^{\prime}}(0,T;\mathrm{L}^{\beta^{\prime}}(\Omega)^{d})$.
In view of (6.6) and a density argument, equation (6.6) may be tested with $\bm{\psi}\in\mathrm{L}^{\infty}(0,T;\mathbf{V})\penalty-10000\cap\mathrm{L}^{\beta}(0,T;\mathbf{V}_{\beta})$, for any $\beta\geq p$, provided that all resulting terms are well defined. This requirement is verified in (6.67)–(6.69). Indeed, there exists a sequence of test functions $\{\bm{\varphi}_{m}\}_{m\in\mathbb{N}}\subset\mathbf{C}^{\infty}(Q_{T})$, with $\operatorname{supp}(\bm{\varphi}_{m})\Subset\Omega\times[0,T)$, such that $\bm{\varphi}_{m}\to\bm{\psi},\ \text{ as }\ m\to\infty$, in $\mathrm{L}^{\infty}(0,T;\mathbf{V})\cap\mathrm{L}^{\beta}(0,T;\mathbf{V}_{\beta})$.
6.2. A priori estimates
As we have seen in Section 4, and as $\beta$ satisfies (4.1), the space of test functions coincides with the space where the solutions to the problem (6.3) are constructed. In particular, the counterparts of (4.10) and (4.11) are also valid here in the same conditions of Lemma 4.1, using in addition the weakly lower semicontinuity of the norm,
$\begin{split}&\sup_{t\in[0,T]}\mathopen{}\mathclose{{\left\{\|\bm{v}_{n}(t)\|^{2}_{2}+\kappa\|\nabla\bm{v}_{n}(t)\|_{2}^{2}}}\right\}+\int_{0}^{T}\mathopen{}\mathclose{{\left(\nu\|\nabla\bm{v}_{n}(t)\|^{p}_{p}+\frac{1}{n}\|\nabla\bm{v}_{n}(t)\|^{\beta}_{\beta}}}\right)\,\mathrm{dt}\\ &\leq C_{1}\mathopen{}\mathclose{{\left(\int_{0}^{T}\|\bm{f}(t)\|^{p^{\prime}}_{p^{\prime}}\,\mathrm{dt}+\|\bm{v}_{0}\|^{2}_{2}+\|\nabla\bm{v}_{0}\|_{2}^{2}}}\right),\end{split}$ (6.8)
and
$\begin{split}&\sup_{t\in[0,T]}\mathopen{}\mathclose{{\left\{\|\nabla\bm{v}_{n}(t)\|^{p}_{p}+\frac{1}{n}\|\nabla\bm{v}_{n}(t)\|^{\beta}_{\beta}}}\right\}+\int_{0}^{T}\mathopen{}\mathclose{{\left(\|{\partial_{t}\bm{v}_{n}}(t)\|_{2}^{2}+\kappa\|\nabla{\partial_{t}\bm{v}_{n}}(t)\|^{2}_{2}}}\right)\,\mathrm{dt}\\ &\leq C_{2}\mathopen{}\mathclose{{\left(\int_{0}^{T}\|\bm{f}(t)\|^{p^{\prime}}_{p^{\prime}}\,\mathrm{dt}+\|\nabla\bm{v}_{0}\|^{p}_{p}+\frac{1}{n}\|\nabla\bm{v}_{0}\|^{\beta}_{\beta}}}\right).\end{split}$ (6.9)
Given their importance to the remainder of this work, the following two estimates are presented as separate claims.
Claim 1.
For any $p\geq 1$, there exists a positive constant $C$ (independent of $n$) such that,
$\int_{0}^{T}\|\bm{v}_{n}(t)\|^{\zeta}_{\zeta}\mathrm{d}t\leq C,\quad\text{ for }\quad\zeta:=\mathopen{}\mathclose{{\left\{\begin{aligned} &\frac{pd}{d-2},\text{ for }d\neq 2,\\ &[1,\infty),\text{ for }d=2.\end{aligned}}}\right$. (6.10)
Proof of Claim 1.
Assume that $d\not=2$ (for $d=2$ is easier).
For $1\leq p\leq 2$, we can use Sobolev’s and Hölder’s inequalities, together with (6.8), so that
$\begin{split}\int_{0}^{T}\|\bm{v}_{n}(t)\|^{\frac{pd}{d-2}}_{\frac{pd}{d-2}}\mathrm{d}t&\leq C\int_{0}^{T}\|\nabla\bm{v}_{n}(t)\|_{2}^{\frac{pd}{d-2}}\mathrm{d}t\\ &\leq C(T)\bigg(\sup_{t\in[0,T]}\mathopen{}\mathclose{{\left\{\|\bm{v}_{n}(t)\|^{2}_{2}+\kappa\|\nabla\bm{v}_{n}(t)\|^{2}_{2}}}\right\}\bigg)^{\frac{pd}{2(d-2)}}\leq C.\end{split}$
If $p>2$, we can use the embedding
$\mathrm{L}^{\infty}(0,T;\mathrm{W}_{0}^{1,2}(\Omega)^{d})\cap\mathrm{L}^{p}(0,T;\mathrm{W}_{0}^{1,p}(\Omega)^{d})\hookrightarrow\mathrm{L}^{\frac{pd}{d-2}}(0,T;\mathrm{L}^{\frac{pd}{d-2}}(\Omega)^{d}),$
together with Sobolev’s, Hölder’s and Korn’s inequalities, and with (6.8), to show that
$\begin{split}\int_{0}^{T}\|\bm{v}_{n}(t)\|^{\frac{pd}{d-2}}_{\frac{pd}{d-2}}\mathrm{d}t&\leq C\int_{0}^{T}\|\nabla\bm{v}_{n}(t)\|_{2}^{\frac{2p}{d-2}}\|\nabla\bm{v}_{n}(t)\|_{p}^{p}\mathrm{d}t,\quad(p>2)\\ &\leq C\bigg(\sup_{t\in[0,T]}\mathopen{}\mathclose{{\left\{\|\bm{v}_{n}(t)\|^{2}_{2}+\kappa\|\nabla\bm{v}_{n}(t)\|^{2}_{2}}}\right\}\bigg)^{\frac{2p}{2(d-2)}}\int_{0}^{T}\|\nabla\bm{v}_{n}(t)\|_{p}^{p}\mathrm{d}t\leq C.\end{split}$
Hence, we obtain the required estimate (6.10). ∎
Claim 2.
For any $p>\frac{2d}{d+2}$, there exist a positive constant $C$ (independent of $n$) such that
$\int_{0}^{T}\|\bm{v}_{n}(t)\otimes\bm{v}_{n}(t)\|^{r_{0}}_{r_{0}}\mathrm{d}t+\int_{0}^{T}\|\mathbf{div}\,(\bm{v}_{n}(t)\otimes\bm{v}_{n}(t))\|^{r_{0}}_{r_{0}}\,\mathrm{dt}\leq C,$ (6.11)
for
$1\leq r_{0}\leq\frac{d}{d-1}.$ (6.12)
Proof of Claim 2.
Assume that $d\not=2$. By Hölder’s inequality and (6.8), one immediately has
$\int_{0}^{T}\|\bm{v}_{n}(t)\otimes\bm{v}_{n}(t)\|^{r_{0}}_{r_{0}}\mathrm{d}t\leq\int_{0}^{T}\|\bm{v}_{n}(t)\|^{2r_{0}}_{2r_{0}}\mathrm{d}t\leq{C(T)\bigg(\sup_{t\in[0,T]}\|\nabla\bm{v}_{n}(t)\|^{2}_{2}\bigg)^{r_{0}}}\leq C,$
for
$1<r_{0}\leq\frac{d}{d-2}.$
In turn, using Hölder’s and Sobolev’s inequalities, together with (6.8), one has
$\begin{split}\int_{0}^{T}\|\mathbf{div}\,(\bm{v}_{n}(t)\otimes\bm{v}_{n}(t))\|^{r_{0}}_{r_{0}}\mathrm{d}t&\leq\int_{0}^{T}\mathopen{}\mathclose{{\left\|\big|\bm{v}_{n}(t)\big|\,\big|\nabla\bm{v}_{n}(t)\big|}}\right\|^{r_{0}}_{r_{0}}\mathrm{d}t\\ &\leq\int_{0}^{T}\|\bm{v}_{n}(t)\|_{\frac{2r_{0}}{2-r_{0}}}^{r_{0}}\|\nabla\bm{v}_{n}(t)\|_{2}^{r_{0}}\mathrm{d}t\\ &\leq C\int_{0}^{T}\|\nabla\bm{v}_{n}(t)\|_{2}^{2r_{0}}\mathrm{d}t,\quad\text{ for }\quad 1\leq r_{0}\leq\frac{d}{d-1}\\ &\leq C(T)\sup_{t\in[0,T]}\|\nabla\bm{v}_{n}(t)\|^{2}_{2r_{0}}\leq C.\end{split}$
Combining the above estimates, we prove (6.11) for $r_{0}\in\Big[1,\frac{d}{d-1}\Big]$.
The case $d=2$ is easier. The first part follows directly from (6.10), while the second part is a straightforward consequence of Sobolev’s inequality. ∎
Using (6.8) and (6.9), together with (6.11), we can apply the Banach-Alaoglu theorem, to pass to the limit along subsequences, still labeled by the same subscript, as follows,
$\bm{v}_{n}$ $\xrightharpoonup[n\to\infty]{\ast}\bm{v},\quad$ $\text{in}\quad\mathrm{L}^{\infty}(0,T;\mathbf{V}),$ (6.13)
$\bm{v}_{n}$ $\xrightharpoonup[n\to\infty]{\ast}\bm{v},\quad$ $\text{in}\quad\mathrm{L}^{\infty}(0,T;\mathbf{V}_{p}),$ (6.14)
$\bm{v}_{n}$ $\xrightharpoonup[n\to\infty]{}\bm{v},\quad$ $\text{in}\quad\mathrm{L}^{p}(0,T;\mathbf{V}_{p}),$ (6.15)
$\mathbf{A}(\bm{v}_{n})$ $\xrightharpoonup[n\to\infty]{}\overline{\bm{\mathcal{A}}},\quad$ $\text{in}\quad\mathrm{L}^{p^{\prime}}(0,T;\mathrm{L}^{p^{\prime}}(\Omega)^{d\times d}),$ (6.16)
$\mathbf{A}(\bm{v}_{n})$ $\xrightharpoonup[n\to\infty]{}\overline{\bm{\mathcal{A}}},\quad$ $\text{in}\quad\mathrm{L}^{p^{\prime}}(0,T;\mathrm{W}^{-1,p^{\prime}}(\Omega)^{d\times d}),$ (6.17)
${\partial_{t}\bm{v}_{n}}$ $\xrightharpoonup[n\to\infty]{}{\partial_{t}\bm{v}},\quad$ $\text{in}\quad\mathrm{L}^{2}(0,T;\mathbf{V}),$ (6.18)
From (6.13), (6.15) and (6.18), it follow assertions (1)-(2) of Definition 3.1. On the other hand, using Korn’s inequality and (6.8), one has
$\int_{0}^{T}\mathopen{}\mathclose{{\left\|\frac{1}{n}|\mathbf{D}(\bm{v}_{n}(t))|^{\beta-2}\mathbf{D}(\bm{v}_{n}(t))}}\right\|_{\beta^{\prime}}^{\beta^{\prime}}\,\mathrm{dt}\leq\frac{1}{n^{\frac{1}{\beta-1}}}\frac{1}{n}\int_{0}^{T}\|\nabla\bm{v}_{n}(t)\|^{\beta}_{\beta}\,\mathrm{dt}\leq\frac{C}{n^{\frac{1}{\beta-1}}},$
which proves that
$\frac{1}{n}|\mathbf{D}(\bm{v}_{n})|^{\beta-2}\mathbf{D}(\bm{v}_{n})\xrightarrow[n\to\infty]{}0,\quad\text{in}\quad\mathrm{L}^{\beta^{\prime}}(0,T;\mathrm{L}^{\beta^{\prime}}(\Omega)^{d\times d}).$ (6.19)
6.3. Pressure recovery and decomposition
In the shear-thinning case ($1<p<2$), and in particular for $p$ satisfying (4.37), the reintroduction and decomposition of the pressure in the weak formulation constitute a structural necessity for establishing an existence result. This procedure enables the precise separation of the inertial, viscous, and convective contributions, as stated in Theorem 5.1. It further compensates for the lack of regularity of the velocity field, by providing the analytical framework required to control nonlinear terms under weaker integrability. In addition, the decomposition plays a key role in compactness and convergence arguments, allowing one to isolate suitable regular pressure components from more singular parts, thereby ensuring strong convergence of the nonlinear expressions in the weak formulation. Without this step, the interplay between the viscous nonlinearity and the limited compactness of the velocity field would obstruct the proof of existence. For that purpose, let us fix
[…]
for
$r_{0}:=\frac{d}{d-1}.$ (6.22)
We emphasize that, in (6.21) and (6.22), we have employed the continuity of $\nabla\Delta^{-1}$ as an operator from $\mathrm{L}^{r_{0}}(\Omega)^{d}$ into $\mathrm{W}^{1,r_{0}}(\Omega)^{d\times d}$, combined with assumption (3.4) and the estimates (6.11) and (6.12). This yields (6.22) with $r_{0}=\min\mathopen{}\mathclose{{\left\{p^{\prime},\,\beta^{\prime},\,\frac{d}{d-1}}}\right\}$. But, due to assumption (6.2), it follows that $r_{0}=\frac{d}{d-1}$. By Theorem 5.1, there exist functions
$\pi_{1,n}$ $\in\mathrm{L}^{p^{\prime}}(0,T;\mathrm{L}^{p^{\prime}}(\Omega)),$
$\pi_{2,n}$ $\in\mathrm{L}^{r_{0}}(0,T;\mathrm{W}^{1,r_{0}}(\Omega)),$
$\pi_{h,n}$ $\in\mathrm{C}_{\mathrm{w}}([0,T];\mathrm{W}^{1,2}(\Omega)),$
with $\Delta\pi_{h,n}=0$ and $\pi_{h,n}(0)=0$ (cf. (5.3)). Moreover, in view of (5.4)-(5.7), (6.8) and (6.20)-(6.21),
$\int_{0}^{T}\|\pi_{1,n}(t)\|_{p^{\prime}}^{p^{\prime}}\,\mathrm{dt}$ $\leq C_{1},$ (6.23)
$\int_{0}^{T}\mathopen{}\mathclose{{\left(\|\pi_{2,n}(t)\|_{r_{0}}^{r_{0}}+\|\nabla\pi_{2,n}(t)\|_{r_{0}}^{r_{0}}}}\right)\,\mathrm{dt}$ $\leq C_{2},$ (6.24)
$\sup_{t\in[0,T]}\|\pi_{h,n}(t)\|_{1,2}$ $\leq C_{3},$ (6.25)
for some positive constants $C_{1}$, $C_{2}$ and $C_{3}$ that do not depend on $n$. In particular, using (5.2) of Theorem 5.1, and (4.5), we deduce
$-\!\int_{0}^{T}\!\!\!\!\int_{\Omega}\big(\bm{v}_{n}{\partial_{t}\bm{\varphi}}+\kappa\nabla\bm{v}_{n}:\nabla{\partial_{t}\bm{\varphi}}\big)\,\mathrm{d}\bm{x}\mathrm{d}t$
[…]
$\quad\quad+\int_{\Omega}\big(\bm{v}_{0,n}\cdot\bm{\varphi}(0)+\kappa\nabla\bm{v}_{0,n}:\nabla\bm{\varphi}(0)\big)\,\mathrm{d}\bm{x}+\!\int_{0}^{T}\!\!\!\!\int_{\Omega}\nabla\pi_{h,n}\cdot{\partial_{t}\bm{\varphi}}\,\mathrm{d}\bm{x}\mathrm{d}t,$ (6.26)
[…]
On the other hand, by (5.3)
$\Delta\pi_{h,n}=0,\quad\mbox{in}\quad\Omega,\qquad\pi_{h,n}(0)=0,$ (6.27)
and so $\pi_{h,n}(t)$ is harmonic in $\Omega$ for any $t\in(0,T)$. By the local regularity theory (see e.g. [27, Theorem 8.24]), and by using (5.4), (5.17)-(5.18), Poincaré’s inequality and (4.5), it follows that for all $t\in(0,T)$ and all $r\in[1,\infty]$,
$\|\pi_{h,n}(t)\|_{2,r}\leq C_{1}\|\pi_{h,n}(t)\|_{2}\leq C_{2}\big(\|\bm{v}_{n}(t)-\bm{v}_{0,n}\|_{2}+\kappa\|\nabla\bm{v}_{n}(t)-\nabla\bm{v}_{0,n}\|_{2}\big),$ (6.28)
for some positive constants $C_{1}$ and $C_{2}$. Hence, by using assumption (3.3), together with (4.7) and (6.13), we have
$\|\pi_{h,n}\|_{\mathrm{L}^{\infty}(0,T;\mathrm{W}^{2,r}_{\mathrm{loc}}(\Omega))}\leq C,\quad\text{ for all }\ r\geq 1.$ (6.29)
Then, by using (6.23)-(6.24) and (6.29), the Banach-Alaoglu theorem gives the existence of functions $\pi_{h}$, $\pi_{1}$ and $\pi_{2}$ such that, for some subsequences, still labeled by the same subscript,
$\pi_{h,n}$ $\xrightharpoonup[n\to\infty]{}\pi_{h},\text{ in }\mathrm{L}^{r}(0,T;\mathrm{W}_{\mathrm{loc}}^{2,r}(\Omega)),\quad\text{ for all }\ r\geq 1,$ (6.30)
$\pi_{1,n}$ $\xrightharpoonup[n\to\infty]{}\pi_{1},\text{ in }\mathrm{L}^{p^{\prime}}(0,T;\mathrm{L}^{p^{\prime}}(\Omega)),$ (6.31)
$\pi_{2,n}$ $\xrightharpoonup[n\to\infty]{}\pi_{2},\text{ in }\mathrm{L}^{r_{0}}(0,T;\mathrm{W}^{1,r_{0}}(\Omega)),$ (6.32)
for $r_{0}$ defined in (6.22). Since the Laplacian $\Delta:\mathrm{W}^{2,r}_{\rm loc}(\Omega)\longrightarrow\mathrm{L}^{r}_{\rm loc}(\Omega)$ is linear and continuous, we can pass (6.27) to the weak limit in $\mathrm{L}^{r}_{\rm loc}(\Omega)$ for almost every $t\in[0,T]$, so that
$\Delta\pi_{h}=0,\quad\mbox{in}\quad\Omega,\qquad\pi_{h}(0)=0.$ (6.33)
6.4. Compactness
Here, we first observe that by using (6.6), together with (3.4), (6.2), (6.8)-(6.9) and (6.10)-(6.11), we can show that
$\big\|\Delta\bm{v}_{n}\big\|_{\mathrm{W}^{1,r_{0}}(0,T;\mathrm{W}^{-1,r_{0}}(\Omega)^{d})}\leq C,$
for $r_{0}$ given by (6.22). As a consequence,
$\big\|\bm{v}_{n}\big\|_{\mathrm{W}^{1,r_{0}}(0,T;\mathrm{W}^{1,r_{0}}(\Omega)^{d})}\leq C,$ (6.34)
Such value of $r_{0}$ implies $r_{0}^{\prime}=d$, and thus we have
$\mathrm{W}_{0}^{1,r_{0}}(\Omega)^{d}\hookrightarrow\mathrm{L}^{2}(\Omega)^{d}\hookrightarrow\mathrm{W}^{-1,r_{0}}(\Omega)^{d},\quad d\leq 4.$
Therefore, from (6.34), we deduce
$\big\|\bm{v}_{n}\big\|_{\mathrm{W}^{1,r_{0}}(0,T;\mathrm{W}^{-1,r_{0}}(\Omega)^{d})}\leq C.$ (6.35)
Observing again the value of $r_{0}$, we can see the following compact and continuous embeddings hold,
$\mathrm{W}_{0}^{1,p}(\Omega)^{d}\hookrightarrow\hookrightarrow\mathrm{L}^{r}(\Omega)^{d}\hookrightarrow\mathrm{W}^{-1,r_{0}}(\Omega)^{d},\qquad 1\leq\ r<\frac{dp}{d-p},\quad p<d,$ (6.36)
or for any finite $r\geq 1$ if $p=d$. Combining this with Lemma 2.2-(1), using the parameters $s=0$, $p=r$, $s_{0}=1$, $p_{0}=r_{0}$, $s_{1}=0$, $p_{1}=p$, we obtain the compact embedding
$\mathrm{W}^{1,r_{0}}(0,T;\mathrm{W}^{-1,r_{0}}(\Omega)^{d})\cap\mathrm{L}^{p}(0,T;\mathrm{W}_{0}^{1,p}(\Omega)^{d})\hookrightarrow\hookrightarrow\mathrm{L}^{r}(0,T;\mathrm{L}^{r}(\Omega)^{d}),$ (6.37)
for every $r$ satisfying (6.36). Hence, for such $r$ and in view of (6.8), (6.35), and (6.37), there exists a subsequence, still labeled by the subscript $n$, such that
$\bm{v}_{n}\xrightarrow[n\to\infty]{}\bm{v},\quad\text{in}\quad\mathrm{L}^{r}(0,T;\mathrm{L}^{r}(\Omega)^{d}).$ (6.38)
Using the Riesz–Fischer theorem, we can extract a further subsequence, still denoted by $n$, for which
$\bm{v}_{n}\xrightarrow[n\to\infty]{}\bm{v},\quad\text{a.e. in }Q_{T},$
and consequently,
$\bm{v}_{n}\otimes\bm{v}_{n}\xrightarrow[n\to\infty]{}\bm{v}\otimes\bm{v},\quad\text{a.e. in }Q_{T}.$ (6.39)
Then, using (6.11) together with (6.39), the same classical result applied in (4.26) yields
$\bm{v}_{n}\otimes\bm{v}_{n}\xrightharpoonup[n\to\infty]{}\bm{v}\otimes\bm{v},\quad\text{in }\mathrm{L}^{r_{0}}(0,T;\mathrm{L}^{r_{0}}(\Omega)^{d\times d})$}, (6.40)
for $r_{0}$ satisfying (6.12), and hence (6.22).
Let us work now on the harmonic pressure term $\{\pi_{h,n}\}_{n\in\mathbb{N}}$, and prove that
$\mathrm{L}^{\infty}(0,T;\mathrm{L}^{2}(\Omega))\cap\{\Delta u(t)=0,\text{ for a.e. }t\}\ \hookrightarrow\hookrightarrow\ {\mathrm{L}^{r}(0,T;\mathrm{L}^{r}_{\rm{loc}}(\Omega))},\quad\text{ for all }\ r>1.$ (6.41)
To prove this, we start by considering an arbitrary but fixed compact set $\Omega^{\prime}\Subset\Omega$, and observe from (6.27) that for each fixed $t$, $\pi_{h,n}(t)$ is harmonic. By standard interior estimates for harmonic functions (see e.g. [27, Theorem 2.10]),
$\|\pi_{h,n}(t)\|_{r,\Omega^{\prime}}\leq C\|\pi_{h,n}(t)\|_{2,\Omega},\quad\text{ for all }\ r\geq 1,$
for some positive constant $C$ independent of $t$. If $\pi_{h,n}\in\mathrm{L}^{\infty}(0,T;\mathrm{L}^{2}(\Omega))$, the right-hand side is uniformly bounded, giving
$\pi_{h,n}\in\mathrm{L}^{r}(0,T;\mathrm{L}^{r}(\Omega^{\prime})),\quad\text{ for all }\ r\geq 1.$
By the Rellich-Kondrachov theorem, after extracting a subsequence,
$\pi_{h,n}(t)\xrightarrow[n\to\infty]{}\pi_{h}(t),\quad\mbox{in}\ \ \mathrm{L}^{r}(\Omega^{\prime}),\quad\text{ for all }\ r>1,$
for a.e. $t$. The uniform bound of $\pi_{h,n}(t)$ in $\mathrm{L}^{r}(\Omega^{\prime})$ allows us to use the Lebesgue dominated convergence theorem to conclude for a further subsequence that
$\pi_{h,n}\xrightarrow[n\to\infty]{}\pi_{h},\quad\text{in}\ \ \mathrm{L}^{r}(0,T;\mathrm{L}^{r}(\Omega^{\prime})).$
Since $\Omega^{\prime}\Subset\Omega$ was chosen arbitrary, this proves that the embedding (6.41) is compact. Hence, for some subsequence still labelled by the subscript $n$, we get from (6.41)
$\pi_{h,n}\xrightarrow[n\to\infty]{}\pi_{h},\quad\text{in}\quad\mathrm{L}^{r}(0,T;\mathrm{L}^{r}_{\rm{loc}}(\Omega)),\quad\text{ for all }\ r>1.$ (6.42)
Let us now fix $r>1$ and compactly contained sets $\Omega^{\prime\prime}\Subset\Omega^{\prime}\Subset\Omega$. For each $n$ and a.e. $t$, the difference
$\varpi_{h,n}(t):=\pi_{h,n}(t)-\pi_{h}(t)$
is harmonic in $\Omega$. By the interior estimate for harmonic functions (see e.g. [27, Theorem 9.11]),
$\|\varpi_{h,n}(t)\|_{2,r,\Omega^{\prime\prime}}\leq C\,\|\varpi_{h,n}(t)\|_{r,\Omega^{\prime}},\quad\text{for a.e.\ }t\in(0,T),$ (6.43)
for some positive constant $C=C(\Omega^{\prime},\Omega^{\prime\prime},r)$. Raising (6.43) to the $r$-th power and integrate over $(0,T)$, we get
$\int_{0}^{T}\|\varpi_{h,n}(t)\|_{2,r,\Omega^{\prime\prime}}^{r}\,\mathrm{dt}\leq C\int_{0}^{T}\|\varpi_{h,n}(t)\|_{r,\Omega^{\prime}}^{r}\,\mathrm{dt}.$
By (6.42), the right-hand side tends to 0 as $n\to\infty$, hence
$\pi_{h,n}\xrightarrow[n\to\infty]{}\pi_{h},\quad\text{in }\mathrm{L}^{r}(0,T;\mathrm{W}^{2,r}(\Omega^{\prime\prime})).$
Since $\Omega^{\prime\prime}\Subset\Omega$ is arbitrary, we conclude
$\pi_{h,n}\xrightarrow[n\to\infty]{}\pi_{h},\quad\text{in }\mathrm{L}^{r}(0,T;\mathrm{W}^{2,r}_{\rm loc}(\Omega)).$ (6.44)
It last to show that $\overline{\bm{\mathcal{A}}}=\mathbf{A}(\bm{v})$.
6.5. Passing to the limit $n\to\infty$
Using the convergence results (6.16), (6.18), (6.19) and (6.40), together with (4.7), (6.20)-(6.21) and (6.30)-(6.32), we can pass to the limit $n\to\infty$ in (6.3) so that the following identity holds:
[…]
for all $\bm{\varphi}\in\mathbf{C}^{\infty}(Q_{T})$ with $\operatorname{supp}(\bm{\varphi})\Subset\Omega\times[0,T)$, where
[…]
Writing the difference between equations (6.45) and (6.3), we have
[…]
for all $\bm{\varphi}\in\mathbf{C}^{\infty}(Q_{T})$ with $\operatorname{supp}(\bm{\varphi})\Subset\Omega\times[0,T)$, and where
[…]
$\varpi_{h,n}$ $:=\pi_{h,n}-\pi_{h},\quad\varpi_{1,n}:=\pi_{1,n}-\pi_{1},\quad\varpi_{2,n}:=\pi_{2,n}-\pi_{2}.$ (6.49)
As a consequence of (6.15), (6.16), (6.19), (6.38) and (6.40), we have
$\bm{v}_{n}-\bm{v}$ $\xrightarrow[n\to\infty]{}\mathbf{0},\quad\text{ in }\quad\mathrm{L}^{r}(0,T;\mathrm{L}^{r}(\Omega)^{d}),$ (6.50)
$\bm{v}_{n}-\bm{v}$ $\xrightharpoonup[n\to\infty]{}\mathbf{0},\quad\text{ in }\quad\mathrm{L}^{p}(0,T;\mathrm{W}_{0}^{1,p}(\Omega)^{d}),$ (6.51)
[…]
for $r_{0}$ and $r$ satisfying (6.12) and (6.36), respectively. On its turn, by (6.31)-(6.32) and (6.44), it is true that
$\varpi_{h,n}$ $\xrightarrow[n\to\infty]{}0,\quad\text{ in }\quad{\mathrm{L}^{r}(0,T;\mathrm{W}^{2,r}_{\rm{loc}}(\Omega)),\quad\text{ for all }\ r>1},$ (6.54)
$\varpi_{1,n}$ $\xrightharpoonup[n\to\infty]{}0,\quad\text{ in }\quad\mathrm{L}^{p^{\prime}}(0,T;\mathrm{L}^{p^{\prime}}(\Omega)),$ (6.55)
$\varpi_{2,n}$ $\xrightharpoonup[n\to\infty]{}0,\quad\text{ in }\quad\mathrm{L}^{r_{0}}(0,T;\mathrm{W}^{1,r_{0}}(\Omega)),$ (6.56)
for $r_{0}$ and $r$ satisfying (6.12) and (6.36), respectively. Moreover, by (6.29), we have
$\varpi_{h,n}$ $\in\mathrm{L}^{\infty}(0,T;\mathrm{W}^{2,r}_{\mathrm{loc}}(\Omega)),\ \text{ for all }\ r\geq 1,$ (6.57)
uniformly in $n$.
Let us introduce the sequence
$\bm{w}_{n}:=\bm{v}_{n}+\nabla\varpi_{h,n},$ (6.58)
and the double sequence
$\bm{w}_{n,j}=\bm{w}_{n}-\bm{w}_{j},\ \text{ where }\ \bm{w}_{j}=\bm{v}_{j}+\nabla\varpi_{h,j},$ (6.59)
and
$\bm{w}_{n,j}(0)=\bm{w}_{n,j}^{0}=\bm{v}_{n}(0)-\bm{v}_{j}(0)+\nabla\varpi_{h,n}(0)-\nabla\varpi_{h,j}(0),$ (6.60)
for $n\geq j$. In view of (6.34), (6.50), (6.54) and (6.56), we have
$\bm{w}_{n,j}$ $\xrightharpoonup[j\to\infty]{}\mathbf{0},\quad\text{ in }\quad\mathrm{L}^{r_{0}}(0,T;\mathrm{W}^{1,r_{0}}(\Omega)^{d\times d}),$ (6.61)
$\bm{w}_{n,j}$ $\xrightarrow[j\to\infty]{}\mathbf{0},\quad\text{in}\quad\ \ \mathrm{L}^{r}(0,T;\mathrm{L}^{r}_{\mathrm{loc}}(\Omega)^{d}),$ (6.62)
for $r_{0}$ and $r$ satisfying (6.12) and (6.36), respectively. From (6.46), we deduce that for $n\geq j$
[…]
for all $\bm{\varphi}\in\mathbf{C}^{\infty}(Q_{T})$ with $\operatorname{supp}(\bm{\varphi})\Subset\Omega\times[0,T)$, and where for each $i\in\{1,2\}$
[…]
$\varpi_{i,n,j}:=\varpi_{i,n}-\varpi_{i,j}.$ (6.65)
Using the definition of the operator $\mathbf{I}-\kappa\Delta$ (see (5.8)), together with (6.49), (6.27), (6.33) and (6.58)–(6.59), and the fact that $\operatorname{div}\nabla^{2}=\nabla\Delta$, we see that
$\begin{split}&\!\int_{0}^{T}\!\!\!\!\int_{\Omega}\Big(\big(\bm{v}_{n}-\bm{v}_{j}+\nabla(\varpi_{h,n}-\varpi_{h,j})\big)\cdot{\partial_{t}\bm{\varphi}}+\kappa\nabla(\bm{v}_{n}-\bm{v}_{j}):\nabla{\partial_{t}\bm{\varphi}}\Big)\,\mathrm{d}\bm{x}\mathrm{d}t\\ &=\!\int_{0}^{T}\!\!\!\!\int_{\Omega}(\mathbf{I}-\kappa\Delta)\bm{w}_{n,j}\cdot{\partial_{t}\bm{\varphi}}\,\mathrm{d}\bm{x}\mathrm{d}t,\end{split}$
and
$\begin{split}&\int_{\Omega}\Big(\big(\bm{v}_{n}(0)-\bm{v}_{j}(0)+\nabla(\varpi_{h,n}(0)-\varpi_{h,j}(0))\big)\cdot\bm{\varphi}(0)+\kappa\nabla(\bm{v}_{n}(0)-\bm{v}_{j}(0)):\nabla{\bm{\varphi}(0)}\Big)\,\mathrm{d}\bm{x}\\ &=\int_{\Omega}(\mathbf{I}-\kappa\Delta)\bm{w}_{n,j}^{0}\cdot\bm{\varphi}(0)\,\mathrm{d}\bm{x}.\end{split}$
Hence (6.63) can be written as follows,
[…]
for all $\bm{\varphi}\in\mathbf{C}^{\infty}(Q_{T})$ with $\operatorname{supp}(\bm{\varphi})\Subset\Omega\times[0,T)$.
We now show that, by testing (6.66) with $\bm{w}_{n,j}$, the right-hand side of (6.66) is well defined. In fact, it suffices to verify that the first and second terms on the right-hand side of (6.66) are well defined when tested with $\bm{w}_{n,j}$.
Let us consider the case $d\neq 2$ (since the case $d=2$ is comparatively easier). Using Sobolev’s embedding, we obtain
$\mathbf{V}{\hookrightarrow}{\mathbf{L}_{\sigma}^{r_{0}^{\prime}}(\Omega)},\quad r_{0}^{\prime}\leq\frac{2d}{d-2}\Leftrightarrow r_{0}\geq\frac{2d}{d+2}.$
From above condition and (6.12), we deduce
$\frac{2d}{d+2}\leq r_{0}\leq\frac{d}{d-1},\quad{d\leq 4}.$ (6.67)
Using integration by parts, Hölder’s and Korn’s inequalities, and still (6.51), (6.52) and (6.55) (to be more specific (6.8), (6.20) and (6.31)), we have
[…]
Using Hölder’s and Sobolev’s inequalities, together with (6.51), (6.53) and (6.56) (to be more specific (6.8), (6.32) and (6.53)), we obtain
[…]
The application of the Minty trick in this case is stated in the following claim.
Claim 3.
$\overline{\bm{\mathcal{A}}}=\mathbf{A}(\bm{v}).$ (6.70)
Proof of Claim 3.
Keep in mind that (6.66) holds for all $\bm{\varphi}\in\mathbf{C}^{\infty}(Q_{T})$ with $\operatorname{supp}(\bm{\varphi})\Subset\Omega\times[0,T)$. We can select our test functions ${\bm{\psi}}\in\mathrm{L}^{\infty}(0,T;\mathbf{V})\cap\mathrm{L}^{p}(0,T;\mathbf{V}_{p})$ (the class of admissible solutions) by using a standard density argument (see (6.67)–(6.69)). This indicates that there is a sequence of test functions $\bm{\varphi}_{m}\in\mathbf{C}^{\infty}(Q_{T})$ with $\operatorname{supp}(\bm{\varphi}_{m})\Subset\Omega\times[0,T)$, such that $\bm{\varphi}_{m}$ converges to $\bm{\psi}$}, in $\mathrm{L}^{\infty}(0,T;\mathbf{V})\cap\mathrm{L}^{p}(0,T;\mathbf{V}_{p})$. Given this information, we test (6.66) with $\bm{w}_{n,j}$ (defined in (6.59)) to get
[…]
where we have used (6.59)–(6.60). Using (6.4) together with (6.60), we deduce that
$J_{1,n,j}\xrightarrow[j\to\infty]{}0.$ (6.72)
By the strong convergence (6.62), we obtain
$\bm{w}_{n,j}\xrightarrow[j\to\infty]{}\mathbf{0},\ \text{ in }\ \mathrm{L}^{r}(0,T;\mathrm{L}^{r}_{\mathrm{loc}}(\Omega)^{d}),$
for $r$ satisfying (6.36). Combining (6.53), (6.56), and the above convergence, we conclude that
$J_{3,n,j}\xrightarrow[j\to\infty]{}0.$ (6.73)
From (6.71), we have
[…]
Using (6.47), (6.58) and (6.64)–(6.65), and performing integration by parts in the resulting second left-hand side term, we get
[…]
According to (6.72) and (6.73),
$J_{1,n,j}+J_{3,n,j}\xrightarrow[j\to\infty]{}0.$
A standard density argument combined with (6.54) lead to
$\mathscr{I}_{n,j}\xrightarrow[j\to\infty]{}0.$
More specifically, we use the strong convergence (see (6.54)) of the harmonic pressure term $\varpi_{h,n}$ on compact subsets of $\Omega\times[0,T)$ to approximate the solution by a sequence of test functions with support compactly embedded in $\Omega\times[0,T)$ that converges to the solution. Therefore, we have
$\limsup_{j\to\infty}\int_{0}^{T}\int_{\Omega}\big(\mathbf{A}(\bm{v}_{n})-\mathbf{A}(\bm{v}_{j})\big):\big(\mathbf{D}(\bm{v}_{n})-\mathbf{D}(\bm{v}_{j})\big)\,\mathrm{d}\bm{x}\mathrm{d}t\leq 0.$ (6.75)
Employing the monotonicity of the nonlinear operator $\mathbf{A}(\cdot)$ (see (2.3)) and (6.75), we find
$\limsup_{j\to\infty}\!\int_{0}^{T}\!\!\!\!\int_{\Omega}\big(\mathbf{A}(\bm{v}_{n})-\mathbf{A}(\bm{v}_{j})\big):\big(\mathbf{D}(\bm{v}_{n})-\mathbf{D}(\bm{v}_{j})\big)\,\mathrm{d}\bm{x}\mathrm{d}t=0.$ (6.76)
Now, set
$\mathcal{F}_{n,j}:=\big(\mathbf{A}(\bm{v}_{n})-\mathbf{A}(\bm{v}_{j})\big):\big(\mathbf{D}(\bm{v}_{n})-\mathbf{D}(\bm{v}_{j})\big).$
Combining (6.76) with Fatou’s lemma, there exists a subsequence $\mathcal{F}_{n,j_{k}}$ such that
$\mathcal{F}_{n,j_{k}}\xrightarrow[k\to\infty]{}0,\qquad\text{for}\quad\lambda^{d+1}\text{-a.e. }(\bm{x},t)\in\Omega\times(0,T).$
Using the strict monotonicity of $\mathbf{A}(\cdot)$, we conclude that
$\mathbf{D}(\bm{v}_{j_{k}})\xrightarrow[k\to\infty]{}\mathbf{D}(\bm{v}_{n}),\qquad\text{for}\quad\lambda^{d+1}\text{-a.e. }(\bm{x},t)\in\Omega\times(0,T).$
Finally, for any $k,\ell$,
$|\mathbf{D}(\bm{v}_{j_{k}})-\mathbf{D}(\bm{v}_{j_{\ell}})|\leq|\mathbf{D}(\bm{v}_{j_{k}})-\mathbf{D}(\bm{v}_{n})|+|\mathbf{D}(\bm{v}_{n})-\mathbf{D}(\bm{v}_{j_{\ell}})|,$
and both terms on the right-hand side converge to zero $\lambda^{d+1}$-a.e. Therefore, $\{\mathbf{D}(\bm{v}_{j})\}_{j\in\mathbb{N}}$ is a Cauchy sequence $\lambda^{d+1}$-almost everywhere. Thus, the limit function exists $\lambda^{d+1}$-a.e. but has to be equal to $\mathbf{D}(\bm{v})$ on account of (6.51), that is,
$\mathbf{D}(\bm{v}_{n})\xrightarrow[n\to\infty]{}\mathbf{D}(\bm{v}),\qquad\text{for}\quad\lambda^{d+1}\text{-a.e. }(\bm{x},t)\in\Omega\times(0,T).$
We address the reader to [20, Eqn. (1.6)] or [24] for more details.
With the help of above facts, we validate the limit process and assure that
$\overline{\bm{\mathcal{A}}}=\mathbf{A}(\bm{v}),\qquad\text{for}\quad\lambda^{d+1}\text{-a.e. }(\bm{x},t)\in\Omega\times(0,T).$
The proof of Claim 3 is therefore over. For a detailed proof of the approach used here, we address the reader to [51, Appendix A] or [14, pp. 324]). ∎
With the above result, the proof of Theorem 3.2 is complete.
7. Uniqueness
The final task in this work is to prove Theorem 3.3, which asserts the uniqueness of weak solutions to the problem (1.1).
Proof.
Let $\bm{v}_{1},\bm{v}_{2}$ be two weak solutions to the problem (1.1) in the sense of Definition 3.1, with the same initial data $\bm{v}_{0}$ and forcing term $\bm{f}$. Define
$\bm{w}:=\bm{v}_{1}-\bm{v}_{2},$
and note that $\mathbf{div}\,\bm{w}=0$ in $Q_{T}$ and $\bm{w}(0)=0$ in $\Omega$. Since $\bm{w}$ is not admissible as a test function in (3.2), we introduce a mollification $\bm{w}^{\varepsilon}:=\rho_{\varepsilon}*\bm{w}$, where the convolution is taken with respect to the time variable. In particular, $\rho_{\varepsilon}(t)$ is supported in $(-\varepsilon,\varepsilon)$. Then $\bm{w}^{\varepsilon}\in\mathbf{C}^{\infty}([0,T];\mathbf{V}_{p})$, $\mathbf{div}\,\bm{w}^{\varepsilon}=0$, $\bm{w}^{\varepsilon}(\cdot,T)=0$ for $\varepsilon$ small enough, and
$\bm{w}^{\varepsilon}\xrightarrow[\varepsilon\to 0]{}\bm{w},\quad\text{in}\quad\mathrm{L}^{2}(0,T;\mathbf{V})\cap\mathrm{L}^{p}(0,T;\mathbf{V}_{p}).$ (7.1)
Using $\bm{w}^{\varepsilon}$ as a test function in (3.2) and subtracting the two weak formulations gives
$\int_{0}^{t}\!\!\int_{\Omega}\partial_{\tau}\bm{w}\cdot\bm{w}^{\varepsilon}\,\mathrm{d}\bm{x}\mathrm{d}\tau+\kappa\int_{0}^{t}\!\!\int_{\Omega}\nabla\partial_{\tau}\bm{w}:\nabla\bm{w}^{\varepsilon}\,\mathrm{d}\bm{x}\mathrm{d}\tau\$
$\ \ +\nu\int_{0}^{t}\!\!\int_{\Omega}\big(\mathbf{A}(\bm{v}_{1})-\mathbf{A}(\bm{v}_{2})\big):\mathbf{D}(\bm{w}^{\varepsilon})\,\mathrm{d}\bm{x}\mathrm{d}\tau$
$=\int_{0}^{t}\!\!\int_{\Omega}(\bm{v}_{1}\otimes\bm{v}_{1}-\bm{v}_{2}\otimes\bm{v}_{2}):\nabla\bm{w}^{\varepsilon}\,\mathrm{d}\bm{x}\mathrm{d}\tau.$
Passing to the limit $\varepsilon\to 0$, using for this purpose (7.1), this yields
$\begin{split}&\frac{1}{2}\|\bm{w}(t)\|_{2}^{2}+\frac{\kappa}{2}\|\nabla\bm{w}(t)\|_{2}^{2}+\nu\int_{0}^{t}\!\!\int_{\Omega}\big(\mathbf{A}(\bm{v}_{1})-\mathbf{A}(\bm{v}_{2})\big):\mathbf{D}(\bm{w})\,\mathrm{d}\bm{x}\mathrm{d}\tau\\ &=\int_{0}^{t}\!\!\int_{\Omega}(\bm{v}_{1}\otimes\bm{v}_{1}-\bm{v}_{2}\otimes\bm{v}_{2}):\mathds{\nabla}\bm{w}\mathrm{d}\bm{x}\mathrm{d}\tau.\end{split}$ (7.2)
Combining the identity
$\bm{v}_{1}\otimes\bm{v}_{1}-\bm{v}_{2}\otimes\bm{v}_{2}=\bm{v}_{1}\otimes\bm{w}+\bm{w}\otimes\bm{v}_{2}$
with the monotonicity of $\mathbf{A}(\cdot)$ (see (2.3)), we obtain
$\begin{split}&\frac{1}{2}\|\bm{w}(t)\|_{2}^{2}+\frac{\kappa}{2}\|\nabla\bm{w}(t)\|_{2}^{2}\leq\int_{0}^{t}\!\!\int_{\Omega}\bm{v}_{1}\otimes\bm{w}:\mathds{\nabla}\bm{w}\mathrm{d}\bm{x}\mathrm{d}\tau+\int_{0}^{t}\!\!\int_{\Omega}\bm{w}\otimes\bm{v}_{2}:\mathds{\nabla}\bm{w}\mathrm{d}\bm{x}\mathrm{d}\tau.\end{split}$ (7.3)
Using integration by parts over $\Omega$, the facts that $\mathbf{div}\,\bm{v}_{1}=0$ in $Q_{T}$ and $\bm{v}_{1}=0$ on $\partial\Omega$ in the sense of trace, the first right-hand side term vanishes. For the second right-hand side term, we use integration by parts, the fact that $\mathbf{div}\,\bm{w}=0$ in $Q_{T}$, Hölder’s and Sobolev’s inequalities, and the fact that $\bm{v}_{2}\in\mathrm{L}^{\infty}(0,T;\mathbf{V})$. After all, we obtain
[…]
for some positive constant $C$. Using this information in (7.3), we get
$\|\nabla\bm{w}(t)\|_{2}^{2}\leq C\int_{0}^{t}\|\nabla\bm{w}(\tau)\|_{2}^{2}\mathrm{d}\tau,$ (7.4)
for a distinct positive constant $C$.
Applying Grönwall’s inequality to (7.4), we obtain
$\|\nabla\bm{w}(t)\|_{2}^{2}\leq\|\nabla\bm{w}(0)\|_{2}^{2}\exp(Ct),\ \text{ for all }\ \ t\in[0,T].$
As $\bm{w}(0)=\bm{0}$, Sobolev’s inequality implies that $\bm{v}_{1}=\bm{v}_{2}$ a.e. in $\Omega$ and for all $t\in[0,T]$ and hence the weak solution to the system (1.1) is unique. ∎
Acknowledgments: A. Kumar acknowledges partial funding by the Austrian Science Fund (FWF), Project Number: P 34681 (10.55776/P34681) and 10.55776/ESP4373225. The work of H. B. de Oliveira was supported by CIDMA (https://ror.org/05pm2mw36) under the Portuguese Foundation for Science and Technology (FCT, https://ror.org/00snfqn58), Grant UID/04106/2025 (https://doi.org/10.54499/UID/04106/2025), and by the Grant no. AP23486218 of the Ministry of Science and Higher Education of the Republic of Kazakhstan. Support for M. T. Mohan’s research received from the National Board of Higher Mathematics (NBHM), Department of Atomic Energy, Government of India (Project No. 02011/13/2025/NBHM(R.P)/R&D II/1137). The first author expresses gratitude to Prof. Dominic Breit, University of Duisburg-Essen for clarifying doubts about his article [14], assistance that benefited this manuscript.
Declarations:
Ethical Approval: Not applicable.
Conflict of interest statement: On behalf of all authors, the corresponding author states that there is no conflict of interest.
Authors’ contributions: All authors have contributed equally.
Funding: Austrian Science Fund (FWF), Project Number: P 34681 (10.55776/P34681) and 10.55776/ESP4373225 (A. Kumar).
Portuguese Foundation for Science and Technology: Grant UID/04106/2025, https://doi.org/10.54499/UID/04106/2025, and Ministry of Science and Higher Education of the Republic of Kazakhstan: Grant AP23486218 (H. B. de Oliveira).
DST-SERB, India, MTR/2021/000066 (M. T. Mohan).
Availability of data and materials: Not applicable.